### v1

# Vision-LSTM: xLSTM as Generic Vision Backbone

Abstract
Transformers are widely used as generic backbones in computer vision, despite initially introduced for natural language processing. Recently, the Long Short-Term Memory (LSTM) has been extended to a scalable and performant architecture – the xLSTM – which overcomes long-standing LSTM limitations via exponential gating and parallelizable matrix memory structure. In this report, we introduce Vision-LSTM (ViL), an adaption of the xLSTM building blocks to computer vision. ViL comprises a stack of xLSTM blocks where odd blocks process the sequence of patch tokens from top to bottom while even blocks go from bottom to top. Experiments show that ViL holds promise to be further deployed as new generic backbone for computer vision architectures.
Project page: https://nx-ai.github.io/vision-lstm/
1 Introduction
Language modeling architectures — such as Transformers [39, 1, 33] or more recently State Space Models [16, 17] such as Mamba [15] — are commonly adapted to the domain of computer vision to make use of their powerful modeling capabilities. However, in natural language processing, an input sentence is typically encoded into tokens that represent words or common subwords [7] via a discrete vocabulary. To encode images into a set of tokens, Vision Transformer [13] (ViT) proposed to group an input image into non-overlapping patches (of e.g. 16x16 pixel), linearly project them into a sequence of so-called patch tokens and add positional information to these tokens. This sequence can then be processed by language modeling architectures.
The Extended Long Short-Term Memory (xLSTM) family [4] was recently introduced as a new architecture for language modeling. It demonstrates the resurgence of LSTM in the LLM era, performing favorably against the likes of Transformers and State Space Models (SSMs). Analogous to existing vision versions of Transformers or State Space Models, e.g., ViT [13] or Vision Mamba [44], which have produced great results in various computer vision tasks [31, 22, 28, 30, 3], we introduce Vision LSTM (ViL) – a generic computer vision backbone that uses xLSTM blocks as its core components. To adjust xLSTM (an autoregressive model) to computer vision (an often non-autoregressive domain), we employ a stack of alternating mLSTM blocks [4] where odd blocks process patches row-wise from top left to bottom right and even blocks go from bottom right to top left. This simple alternating design allows ViL to efficiently process non-sequential inputs, such as images, without introducing additional computations.
Similar to vision adaptions of SSMs [23, 44], ViL exhibits linear computational and memory complexity w.r.t. sequence length which makes it appealing for tasks that benefit from high-resolution images such as medical imaging [8, 18, 38, 41], segmentation [22, 9], or physics simulations [5, 27, 6, 2]. In contrast, ViT’s computational complexity scales quadratically due to the self-attention mechanism, rendering them costly to apply to high-resolution tasks.
2 Method
Vision-LSTM (ViL) is a generic backbone for computer vision tasks, which is residually built from xLSTM blocks, as visualized in Figure 1. Following ViT [13], ViL first splits an image into non-overlapping patches via a shared linear projection, then adds learnable positional embeddings to each patch token. At the core of ViL are alternating mLSTM blocks, which are fully parallelizable and equipped with a matrix memory and a covariance update rule. Odd mLSTM blocks process patch tokens from top left to bottom right while even blocks go from bottom right to top left.
Figure 1: Schematic overview of Vision-LSTM (ViL). Following ViT [13], an input image is split into patches and linearly projected. Then, a learnable vector is added per position to the patches, producing a sequence of patch tokens. This sequence is then processed by alternating mLSTM blocks where even blocks flip the sequence before and after the mLSTM layer. For classification, ViL averages the first and the last patch followed by a linear classification head.
3 Experiments
We conduct experiments on ImageNet-1K [12], which contains 1.3M training images and 50K validation images where each image belongs to one of 1000 classes. Our comparison focuses on models that use a sequence modeling backbone and are roughly comparable in parameter counts.
We train ViL models for 800 epochs (tiny, tiny+) or 400 epochs (small, small+, base) on 224x224 resolution with a learning rate of 1e-3 using a cosine decay schedule. Detailed hyperparameters can be found in Appendix 5.
To facilitate a fair comparison against Vision Mamba (Vim) [44], we add additional blocks to our models to match parameter counts for the tiny and small variants (denoted as ViL-T+ and ViL-S+). Note that ViL requires significantly less compute than Vim since ViL traverses the sequence in an alternating way, whereas Vim traverses the sequence twice per block. This holds despite Vim using optimized CUDA kernels, which currently do not exist for mLSTM, and would speed-up ViL even more. We compare runtimes in Appendix A.1 where ViL is up to 69% faster than Vim.
As ViTs are well established in the vision community, they underwent multiple optimization cycles over the years [13, 34, 36, 35, 37, 19]. As this work is the first to apply xLSTM to computer vision, we do not expected to outperform years of hyperparameter tuning of ViTs in all cases. Even so, the results in Table 1 show that ViL shows favorable results to heavily optimized ViT protocols (DeiT, DeiT-II, DeiT-III) on smaller scales where only DeiT-III-S that was trained twice as long as ViL-S performs slightly better. On the “base” scale, ViL outperforms the initial ViT [13] models and achieves comparable results to DeiT [34]. Note that due to high costs of training models at this scale, the hyperparameters for ViL-B are far from optimal. For reference, training ViL-B takes roughly 600 A100 GPU-hours or 19 hours on 32 A100 GPUs.
The performance can be further improved by fine-tuning models in a “long-sequence fine-tuning” setting [44], which fine-tunes the models for 30 epochs with an increased sequence length of 729 by using 50% overlap between consecutive patch tokens.
ViL also demonstrates competitive performance with CNN-based models such as ConvNeXt [24], despite not leveraging the inductive bias inherent to convolutions.
Method #Params ACC
DeiT-T [34] 6M 72.2
DeiT-II-T [36] 6M 73.5
DeiT-III-T (reimpl.) 6M 76.2
Vim-T [44] 7M 76.1
ViL-T 6M 77.3
ViL-T+ 7M 78.1
DeiT-S [34] 22M 79.8
DeiT-II-S [36] 22M 80.7
DeiT-III-S [37] 22M 81.4
Vim-S [44] 26M 80.5
ViL-S 23M 80.7
ViL-S+ 26M 80.9
ViT-B384 [13] 86M 77.9
DeiT-B [34] 86M 81.8
DeiT-II-B [36] 86M 82.7
DeiT-III-B [37] 86M 83.7
ViL-B 88M 81.6*
Method #Params ACC
Long-sequence fine-tuning
DeiT-III-T (reimpl.) 6M 78.3
Vim-T [44] 7M 78.3
ViL-T+ 7M 80.0
DeiT-S [34] 22M 81.4
DeiT-III-S [37] 22M 83.2
Vim-S [44] 26M 81.6
ViL-S+ 26M 82.2
DeiT-B [34] 86M 83.2
DeiT-III-B [37] 86M 84.5
ViL-B 88M 82.7*
CNN-based models
ConvNeXt-T [24] 30M 82.2
ConvNeXt-S [24] 50M 83.1
ConvNeXt-B [24] 89M 83.8
SSM-based models
S4ND-Conv-T [26] 30M 82.2
S4ND-ViT-B [26] 89M 80.4
Table 1: ImageNet-1K classification accuracy. Entries in gray are trained trained for 400 epochs more than ViL. “Long-sequence fine-tuning” fine-tunes models with 50% overlapping patches for 30 epochs. We train DeiT models in the long-sequence fine-tuning setting as DeiT-II has no public checkpoints and DeiT-III only published checkpoints with longer training durations. *due to the large costs of training ViL-B, hyperparameters are far from optimal.
Block Design.
We study different ways to design ViL blocks in Table 2. The plain, uni-directional xLSTM block fails to reach competitive performances since the autoregressive nature of xLSTM is not suited for image classification. Traversing blocks in a bi-directional manner, i.e., introducing a second mLSTM layer in each block that traverses the sequence backwards (akin to Vim [44]), improves performance, but also requires more parameters and FLOPS. Sharing the parameters of the forward and backward mLSTM makes the model more parameter efficient, but still requires more compute and overloads these parameters which leads to performance drops. Using alternating blocks enhances performance while remaining compute and parameter efficient. We also explore quad-directional designs (similar to [23]), which refers to traversing the sequence row-wise (in both directions) and additionally column-wise (in both directions). Bi-directional traverses the sequence only row-wise (in both directions). Figure 2 visualizes the different traversal paths.
Figure 2: Visualization of traversal paths. Bi-directional uses the first two traversal paths. Quad-directional uses all of the above. Squares represent individual patch tokens.
Due to the increased cost of bi-directional and quad-directional blocks, this study was conducted in a heavily reduced setting. We train on a subset of ImageNet-1K that contains only samples from 100 classes for 400 epochs at 128x128 resolution. This was particularly necessary because our quad-directional implementation is not compatible with torch.compile (a generic speed optimization method from PyTorch [29]), which leads to longer runtimes, as indicated last column of Table 2. Due to this technical limitation, we choose the alternating bi-directional block as our core design.
Direction Alternating Param. Sharing ACC #Param FLOPS Runtime
Uni-dir. (xLSTM) 72.8 6.1M 1.0x 1.0x
Bi-dir. 74.8 6.5M 1.4x 1.9x
Bi-dir. 73.3 6.1M 1.4x 1.9x
Bi-dir. (ViL) 74.8 6.1M 1.0x 1.0x
Quad-dir. 76.3 7.3M 2.3x 7.5x
Quad-dir. 72.8 6.1M 2.3x 7.5x
Quad-dir. 75.6 6.1M 1.0x 1.9x
Table 2: Traversing the sequence of patch tokens in multiple directions yields better performance than the plain, uni-directional xLSTM block. However, using multiple directions per block also requires more compute and parameters. Sharing parameters between directions matches #Params but overloads them, leading to performance drops. Using multi-directional designs but alternating between them in consecutive blocks, matches #Params and FLOPS while achieving good performances. While quad-directional designs yields the best performances, they comes with drastically increased runtime. As our current quad-directional implementation is not compatible with torch.compile, it suffers from increased runtime even when alternating between directions. Since quad-directional designs are compute intensive, this study was conducted by training a ViL-T on an ImageNet-1K subset of 100 classes using resolution 128.
3.1 Classification Design
In order to perform classification with a ViT, the sequence of tokens is pooled into a single token, which is then used as input to a classification head. The most common methods to do this pooling are (i) adding a learnable [CLS] token at the start of the sequence or (ii) averaging all patch tokens to produce an [AVG] token. Whether to use the [CLS] or [AVG] token is typically a hyperparameter, where both variants achieve comparable performances. On the contrary, autoregressive models often require specialized classification designs. For example, Vim [44] requires the [CLS] token to be in the middle of the sequence, suffering heavy performance losses if other classification designs, e.g., an [AVG] token or two [CLS] tokens at start and end of the sequence, are employed. Due to its autoregressive nature, we explore different classification designs for ViL in Table 3. [AVG] averages all patch tokens, “Middle Patch” uses the middle patch token, “Middle [CLS]” uses a [CLS] token in the middle of the sequence, “Bilateral [AVG]” uses the average of the first and last patch token. We find ViL to be relatively robust the classification design where all performances are within 0.6%. We choose “Bilateral [AVG]” over “Middle [CLS]”, as ImageNet-1K is known to have a center bias, i.e., objects are commonly located in the middle of the picture. By using “Bilateral [AVG]” we avoid exploiting this bias, which makes our model more general.
Arch ACC
[AVG] 75.4
Middle Patch 75.9
Middle [CLS] 75.3
Bilateral [AVG] 75.9
Bilateral Concat 76.4
Table 3: How to pool the sequence of patch tokens into a single token for classification? ViL is relatively robust to the classification design. Using the middle patch token and averaging the first and last token (“Bilateral [AVG]”) work similarly well. The good performance of middle patch could be caused by the bias in ImageNet-1K that objects are typically in the center of the image. We therefore choose “Bilateral [AVG]” to avoid exploiting this bias. “Bilateral Concat” concatenates the first and last patch token which introduces additional parameters in the classification head. To save compute, we conduct this study by training a ViL-T on ImageNet-1K for 400 epochs in 192x192 resolution followed by fine-tuning for 20 epochs on 224x224 resolution.
To remain comparable to previous architectures that use a single token as input to the classification head, we average the first and the last patch. To achieve best performances, we suggest to concatenate the two tokens (“Bilateral Concat”) instead of averaging them. This is similar to common practices in self-supervised vision transformers, such as DINOv2 [28], which are trained with two objectives that are attached at the [CLS] and [AVG] token respectively and therefore benefit from concatenating the representation of [CLS] and [AVG] token. This directions has also been explored for vision SSM models [40] where multiple [CLS] tokens are spread out over the sequence and then used as input to a classifier. Similar directions could also improve the performance of ViL.
4 Conclusion
Motivated by the success of xLSTM in language modeling, we introduced ViL, an adaption of the xLSTM architecture to vision tasks. ViL processes a sequence of patch tokens in alternating fashion. Odd blocks process image patches row-wise from top left to bottom right and even blocks go from bottom right to top left. Our new architecture outperforms SSM-based vision architectures and also optimized ViT models on ImageNet-1K classification. Remarkably, ViL is able to outperform ViT training pipelines, that resulted from years of hyperparameter tuning and transformer improvements, in fair comparisons.
In the future, we see potential in applying ViL when high-resolution images are needed for optimal performance, such as semantic segmentation or medical imaging. In these settings, transformers suffer from high computational costs due to the quadratic complexity of self-attention, which ViL does not have due to its linear complexity. Additionally, improving pre-training schemes (e.g. via self-supervised learning), exploring better hyperparameter settings or transferring techniques from transformers (e.g. LayerScale [35]) are promising directions of ViL.
Acknowledgments
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, MeluXina at LuxProvide, Luxembourg, Leonardo at CINECA, Italy and LUMI at CSC, Finland.
The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects Medical Cognitive Computing Center (MC3), INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), AIRI FG 9-N (FWF-36284, FWF-36235), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.
Appendix A Extended Results
A.1 Runtime Comparison of ViL vs Vim
We compare the time it takes to train ViL and Vim [44] for a single ImageNet-1K epoch in Table 4. We follow the scaling procedure of ViTs, using 192 (T), 384 (S), 768 (B), 1024 (L) as latent dimension where the large scale additionally doubles the number of blocks.
Optimization (T)iny (S)mall (B)ase (L)arge
Vim [44] custom CUDA kernel 7.3h 14.0h 28.2h 76.4h
ViL torch.compile 5.0h 8.7h 16.6h 45.1h
Speedup of ViL compared to Vim 45% 61% 69% 69%
Table 4: Runtime comparisons between Vim [44] and ViL. ViL is up to 69% faster despite the current lack of a custom CUDA kernel. As mLSTM (and ViL) can be parallelized analogous to FlashAttention [11, 10] via custom hardware optimizations, ViL will become even faster in the future. Runtimes denote the training time for 10 ImageNet-1K epochs and are extrapolated from short benchmark runs on a single A100-80GB-PCIe using float16 precision and 224x224 images.
A.2 Impact of Longer Training
We investigate training for a longer duration on ViL-T+ and train it for 400 and 800 epochs following the hyperparameters in Table 5. After 400 epochs, ViL-T+ achieves 77.2% accuracy and after doubling the training duration the model achieves 78.1% accuracy.
Our implementation of DeiT-III-T achieves 75.6% after 400 epochs and 76.2% with 800 epochs using the hyperparameters in Table 6.
Appendix B Implementation Details
B.1 Hardware
We train models on a mix of custom hardware servers (mainly A100 and A40 GPUs) and public research clusters equiped with either 8xA100 or 4xA100 nodes.
We estimate the total number of A100 GPU-hours used for this project to be 25K hours. This estimate includes everything from initial exploration, method development, analysis and evaluations.
B.2 FLOPS Calculation
We calculate the FLOPS from Table 2 with the fvcore library [^1] https://github.com/facebookresearch/fvcore. As fvcore does not support all operations of ViL, the FLOPS are not 100% accurate, but nevertheless serve as a good reverence point to compare the relative computational effort between block designs.
B.3 ViL Hyperparameters
Table 5 shows detailed hyperparameters used to train ViL models. In the “long-sequence fine-tuning” setting, we follow Vim [44] and fine-tune for 30 epochs with a learning rate of 1e-5, a weight decay of 1e-8 using an overlap of 50% (8 pixels) between consecutive patches. This increases the sequence length from 196 to 729. Other hyperparameters follow Table 5.
Parameter Value
Epochs 800 (T), 400 (S/B)
Batch size 1024 (S), 2048 (T/B)
Model
   Patch size 16x16
   Latent dimension 192 (T), 384 (S), 768 (B)
   Depth 24 (T, S, B), 29 (T+), 26 (S+)
Stochastic depth [20]
   Peak rate 0 (T), 0.05 (S), 0.2 (B)
   Layer-wise Decay
Optimizer AdamW [25, 21]
   Base Learning rate 1e-3
   Linear LR Scaling Divisor [14] 1024
   Weight decay 0.05
   Momentum $\beta_{1}=0.9,\beta_{2}=0.999$
   Gradient Norm Clip 1.0
Learning rate schedule linear warmup $\rightarrow$ cosine decay
   Warmup epochs 5
   End LR 1e-6
Label smoothing [32]
Train Data Augmentation
   RandomResizedCrop 224
    Scale [0.08, 1.0]
    Interpolation bicubic
   RandomHorizontalFlip $p=0.5$
   3-Augment [37]
    Gaussian Blur $\sigma$ [0.1, 2.0]
    ColorJitter [0.3, 0.3, 0.3, 0.0]
   Normalize ImageNet-1K statistics
   Mixup [43] $\alpha$ 0.8
   Cutmix [42] $\alpha$ 1.0
Test Data Augmentation
   Resize 224
    Interpolation bicubic
   CenterCrop 224
   Normalize ImageNet-1K statistics
Table 5: Hyperparameters for training ViL on ImageNet-1K, inspired by DeiT-III [37].
B.4 DeiT-III Reimplementation Hyperparameters
Table 5 shows detailed hyperparameters used to train DeiT-III-T (reimpl.) from Table 1.
Parameter Value
Epochs 800 $\rightarrow$ 20
Batch size 2048
Stochastic depth [20]
Layerscale [35] 1e-4
Optimizer AdamW [25, 21]
   Base Learning rate 1e-3
   Linear LR Scaling Divisor [14] 1024
   Weight decay 0.05
   Momentum $\beta_{1}=0.9,\beta_{2}=0.999$
   Gradient Norm Clip
Learning rate schedule linear warmup $\rightarrow$ cosine decay
   Warmup epochs 5
   End LR 1e-6
Label smoothing [32]
Train Data Augmentation
   RandomResizedCrop 192 $\rightarrow$ 224
    Scale [0.08, 1.0]
    Interpolation bicubic
   RandomHorizontalFlip $p=0.5$
   3-Augment [37]
    Gaussian Blur $\sigma$ [0.1, 2.0]
    ColorJitter [0.3, 0.3, 0.3, 0.0]
   Normalize ImageNet-1K statistics
   Mixup [43] $\alpha$ 0.8
   Cutmix [42] $\alpha$ 1.0
Test Data Augmentation
   Resize 192 $\rightarrow$ 224
    Interpolation bicubic
   CenterCrop 192 $\rightarrow$ 224
   Normalize ImageNet-1K statistics
Table 6: Hyperparameters for training our reimplementation of DeiT-III-T [37] on ImageNet-1K. The most significant change is that we reduce the learning rate from 3e-3 to 1e-3 as we found this to greatly improve performance. We make minor changes to the protocol such as using AdamW or no gradient clipping as models were stable without it. We follow the best setting from DeiT-III [37] and pre-train on 192 resolution followed by a short fine-tuning on 224 resolution (indicated by $\rightarrow$).

### v2

@@
 This simple alternating design allows ViL to efficiently process non-sequential inputs, such as images, without introducing additional computations.
-Similar to vision adaptions of SSMs [23, 44], ViL exhibits linear computational and memory complexity w.r.t.
+Similar to vision adaptions of SSMs [40, 67, 58], ViL can exhibit linear computational and memory complexity w.r.t.
 sequence length which makes it appealing for tasks that benefit from high-resolution images such as medical imaging [9, 26, 55, 62], segmentation [37, 10], or physics simulations [6, 43, 7, 2].
@@
 Following ViT [18], ViL first splits an image into non-overlapping patches via a shared linear projection, then adds learnable positional embeddings to each patch token.
-At the core of ViL are alternating mLSTM blocks, which are fully parallelizable and equipped with a matrix memory and a covariance update rule.
+At the core of ViL are alternating mLSTM blocks, which are fully parallelizable and equipped with a matrix memory combined with a covariance update rule.
 Odd mLSTM blocks process patch tokens from top left to bottom right while even blocks go from bottom right to top left.
@@
 This sequence is then processed by alternating mLSTM blocks where even blocks flip the sequence before and after the mLSTM layer.
-For classification, ViL averages the first and the last patch followed by a linear classification head.
+For classification, ViL uses the concatenation of the first and the last patch as input to a linear classification head.
 3 Experiments
-We conduct experiments on ImageNet-1K [12], which contains 1.3M training images and 50K validation images where each image belongs to one of 1000 classes.
-Our comparison focuses on models that use a sequence modeling backbone and are roughly comparable in parameter counts.
-We train ViL models for 800 epochs (tiny, tiny+) or 400 epochs (small, small+, base) on 224x224 resolution with a learning rate of 1e-3 using a cosine decay schedule.
-Detailed hyperparameters can be found in Appendix 5.
-To facilitate a fair comparison against Vision Mamba (Vim) [44], we add additional blocks to our models to match parameter counts for the tiny and small variants (denoted as ViL-T+ and ViL-S+).
-Note that ViL requires significantly less compute than Vim since ViL traverses the sequence in an alternating way, whereas Vim traverses the sequence twice per block.
-This holds despite Vim using optimized CUDA kernels, which currently do not exist for mLSTM, and would speed-up ViL even more.
-We compare runtimes in Appendix A.1 where ViL is up to 69% faster than Vim.
+We pre-train models on ImageNet-1K [17], which contains 1.3M training images and 50K validation images where each image belongs to one of 1000 classes.
+ViL models are trained for 800 epochs (tiny) or 400 epochs (small, base) on 192x192 resolution with a learning rate of 1e-3 using a cosine decay schedule.
+Afterwards, the model is fine-tuned on 224x224 resolution for 20 epochs using a learning rate of 1e-5.
+Detailed hyperparameters can be found in Appendix Table 12.
+As ViL is an isotropic architecture (like ViT [18]), we compare it against other isotropic architectures and leave exploration of hierarchical ViL architectures to future work.
+Figure 2: Performance overview of ImageNet-1K pre-trained models in relation to pre-training compute. ViL shows strong performances across classification and semantic segmentation tasks.
 As ViTs are well established in the vision community, they underwent multiple optimization cycles over the years [18, 51, 53, 52, 54, 27].
-As this work is the first to apply xLSTM to computer vision, we do not expected to outperform years of hyperparameter tuning of ViTs in all cases.
-Even so, the results in Table 1 show that ViL shows favorable results to heavily optimized ViT protocols (DeiT, DeiT-II, DeiT-III) on smaller scales where only DeiT-III-S that was trained twice as long as ViL-S performs slightly better.
-On the “base” scale, ViL outperforms the initial ViT [13] models and achieves comparable results to DeiT [34].
-Note that due to high costs of training models at this scale, the hyperparameters for ViL-B are far from optimal.
-For reference, training ViL-B takes roughly 600 A100 GPU-hours or 19 hours on 32 A100 GPUs.
-The performance can be further improved by fine-tuning models in a “long-sequence fine-tuning” setting [44], which fine-tunes the models for 30 epochs with an increased sequence length of 729 by using 50% overlap between consecutive patch tokens.
-ViL also demonstrates competitive performance with CNN-based models such as ConvNeXt [24], despite not leveraging the inductive bias inherent to convolutions.
-Method #Params ACC
-DeiT-T [34] 6M 72.2
-DeiT-II-T [36] 6M 73.5
-DeiT-III-T (reimpl.) 6M 76.2
-Vim-T [44] 7M 76.1
-ViL-T 6M 77.3
-ViL-T+ 7M 78.1
-DeiT-S [34] 22M 79.8
-DeiT-II-S [36] 22M 80.7
-DeiT-III-S [37] 22M 81.4
-Vim-S [44] 26M 80.5
-ViL-S 23M 80.7
-ViL-S+ 26M 80.9
-ViT-B384 [13] 86M 77.9
-DeiT-B [34] 86M 81.8
-DeiT-II-B [36] 86M 82.7
-DeiT-III-B [37] 86M 83.7
-ViL-B 88M 81.6*
-Method #Params ACC
-Long-sequence fine-tuning
-DeiT-III-T (reimpl.) 6M 78.3
-Vim-T [44] 7M 78.3
-ViL-T+ 7M 80.0
-DeiT-S [34] 22M 81.4
-DeiT-III-S [37] 22M 83.2
-Vim-S [44] 26M 81.6
-ViL-S+ 26M 82.2
-DeiT-B [34] 86M 83.2
-DeiT-III-B [37] 86M 84.5
-ViL-B 88M 82.7*
-CNN-based models
-ConvNeXt-T [24] 30M 82.2
-ConvNeXt-S [24] 50M 83.1
-ConvNeXt-B [24] 89M 83.8
-SSM-based models
-S4ND-Conv-T [26] 30M 82.2
-S4ND-ViT-B [26] 89M 80.4
-Table 1: ImageNet-1K classification accuracy.
-Entries in gray are trained trained for 400 epochs more than ViL.
-“Long-sequence fine-tuning” fine-tunes models with 50% overlapping patches for 30 epochs.
-We train DeiT models in the long-sequence fine-tuning setting as DeiT-II has no public checkpoints and DeiT-III only published checkpoints with longer training durations.
-*due to the large costs of training ViL-B, hyperparameters are far from optimal.
-Block Design.
+Since this work is the first to apply xLSTM to computer vision, we do not expected to outperform years of hyperparameter tuning of ViTs in all cases.
+Even so, Figure 2 shows an overview of performance metrics in relation to total pre-training compute where ViL performs favorably against heavily optimized transformer protocols (DeiT, DeiT-III) and Vision Mamba (Vim) [67] on ImageNet-1K [17] classification, ADE20K [66] semantic segmentation and VTAB-1K [64] classification.
+Detailed results are presented in the following sections.
+3.1 ImageNet-1K Pre-training
+Table 1 relates parameter counts and FLOPS to classification accuracy after pre-training on ImageNet-1K.
+ViL outperforms heavily optimized ViT protocols and other backbones on the tiny and small scale.
+While ViL does not outperform all other models on the base scale, evaluations on downstream tasks (as shown later in Table 3 and Table 2) show that ViL-B still learns strong features.
+Additionally, training a ViL-B for 400 epochs takes roughly 600 A100 GPU-hours or 19 hours on 32 A100 GPUs.
+Therefore, the hyperparameter configuration is most likely not optimal as we were not able to extensively optimize hyperparameters for this scale.
+Note that training ViL is still quite expensive to train as there does not exist an optimized hardware implementation yet, which does exist for other models like ViT [15, 14] or Vim [23, 67].
+Nevertheless, the efficient alternating block design of ViL makes it already faster than Vim (up to 69% as shown in Appendix B.1) while only using torch.compile, a generic speed optimization of the PyTorch [45] framework.
+Model Epochs #Params FLOPS IN-1K
+DeiT-T [51] 300 6M 1.3G 72.2
+DeiT-II-T [53] 400 6M 1.3G 73.5
+DeiT-III-T (reimpl.) 800+20 6M 1.3G 76.2
+VRWKV-T [19] 300 6M 1.2G 75.1
+Vim-T [67] 300 7M 1.5G 76.1
+Mamba®-T [58] 280+20 9M - 77.4
+ViL-T 800+20 6M 1.5G 78.3
+DeiT-S [51] 300 22M 4.6G 79.8
+DeiT-II-S [53] 400 22M 4.6G 80.7
+DeiT-III-S [54] 800+20 22M 4.6G 81.4
+ConvNeXt-S [41] (iso.) 300 22M 4.3G 79.7
+VRWKV-S [19] 300 24M 4.6G 80.1
+Vim-S [67] 300 26M 5.3G 80.5
+Mamba®-S [58] 280+20 28M - 81.1
+ViL-S 400+20 23M 5.1G 81.5
+DeiT-B [51] 300 86M 17.6G 81.8
+DeiT-II-B [53] 400 86M 17.6G 82.7
+DeiT-III-B [54] 800+20 86M 17.6G 83.7
+ConvNeXt-B [41] (iso.) 300 87M 16.9G 82.0
+VRWKV-B [19] 300 94M 18.2G 82.0
+Mamba®-B [58] 280+20 99M - 82.9
+ViL-B 400+5 89M 18.6G 82.4
+Table 1: ImageNet-1K pre-training accuracy.
+All models use a patch size of 16x16 with 224x224 resolution at most.
+Models with “+” in their “Epochs” column pre-train on lower resolution followed by fine-tuning on 224x224 resolution for some epochs.
+ViL performs favorably against against an isotropic convolutional architecture (ConvNeXt) and vision adaptions of transformers (DeiT), RWKV [47] (VRWKV) and Mamba [23] (Vim, Mamba®).
+3.2 ADE20K Semantic Segmentation
+Table 2 shows results for transferring ImageNet-1K pre-trained models to ADE20K [66] semantic segmentation using UperNet [61].
+Also here, ViL shows strong performances across the board, even outperforming DeiT-III-B despite the lower ImageNet-1K accuracy of ViL-B.
+For this benchmark, we train on 512x512 images and inference could be conducted on the original images (without resizing), which results in even larger images.
+Consequently, the number of patch tokens can become quite large.
+As sequence length increases, ViL would ideally use a recurrent or chunkwise formulation of the mLSTM layer.
+Shortly summarized, the mLSTM can operate in parallel, recurrent or chunked modes, each with distinct FLOPS and runtime characteristics.
+Given a sequence length $L$ and hidden dimension $D$, the trade-off between modes can be formulated as $\mathcal{O}(LCD+LD^{2})$ where $C$ denotes the chunksize of the chunked form.
+$C=1$ (recurrent mode) yields complexity $\mathcal{O}(LD^{2})$ and $C=L$ (parallel mode) gives $\mathcal{O}(L^{2}D)$ or $\mathcal{O}(L\frac{L}{2}D)$ when omitting the causally masked operations.
+Due to this trade-off and the current lack of an optimized hardware implementation for the mLSTM, it is difficult to create runtime or FLOPS estimates that are of practical relevance in this setting and therefore leave it to future work.
+Single-scale Multi-scale
+Model #Params mIoU ACC mIoU ACC
+DeiT-T [51] 10M 38.1 78.2 40.3 79.9
+DeiT-III-T [54] 10M 39.8 79.2 42.2 80.7
+Vim-T [67] 13M 41.0 - - -
+ViL-T 11M 41.2 80.2 43.1 81.3
+DeiT-S [51] 41M 43.1 80.7 45.2 81.8
+DeiT-III-S [54] 41M 45.2 81.5 46.3 82.3
+Vim-S [67] 46M 44.9 - - -
+Mamba®-S [58] 56M 45.3 - - -
+ViL-S 42M 46.3 82.0 47.9 82.9
+DeiT-B [51] 113M 45.8 82.1 47.0 82.9
+DeiT-III-B [54] 113M 47.5 82.6 49.0 83.3
+Mamba®-B [58] 132M 47.7 - - -
+ViL-B 115M 48.6 82.8 49.6 83.3
+Table 2: Semantic segmentation results on ADE20K [66] using UperNet [61].
+We report mean intersection over union (mIoU) and pixelwise accuracy (ACC) for single- and multi-scale evaluation.
+Models are trained for 160K updates with a batchsize of 16 on 512x512 resolution.
+Detailed hyperparameters are listed in Appendix Table 14.
+3.3 VTAB-1K Transfer Classification
+Table 3 shows transfer classification results for ImageNet-1K pre-trained models on the VTAB-1K [64] benchmark.
+VTAB-1K consists of 19 datasets split into 7 natural datasets (such as CIFAR100 [38], Caltech101 [20] or SUN397 [60]), 4 specialized datasets (medical imaging [57, 34] and remote sensing [28, 11]) and 8 structured datasets (with tasks such as object counting [33] or binned depth estimation [21]).
+We follow common practices and tune the learning rate per model and dataset on the validation set followed by training with the best learning rate on the union of train and validation set.
+The performance metric is the average testset accuracy over 5 seeds.
+ViL shows strong transfer classification performance outperforming all other models on the average over all 19 datasets.
+ViL performs particularly well on the structured datasets where ViL-B outperforms DeiT-III-B despite ViL-B having lower ImageNet-1K accuracy.
+Model #Params FLOPS Natural Specialized Structured Average
+DeiT-T [51] 6M 1.3G 69.2 82.0 53.3 65.2
+DeiT-III-T (reimpl.) 6M 1.3G 71.9 82.6 55.2 67.1
+Vim-T [67] 7M 1.5G 68.0 80.7 47.1 61.9
+ViL-T 6M 1.5G 73.6 83.4 56.1 68.3
+DeiT-S [51] 22M 4.6G 73.3 83.8 53.2 67.1
+DeiT-III-S [54] 22M 4.6G 75.0 83.2 52.3 67.2
+Vim-S [67] 26M 5.3G 69.6 81.7 49.4 63.6
+ViL-S 23M 5.1G 75.3 84.3 58.3 70.0
+DeiT-B [51] 86M 17.6G 76.5 85.2 55.7 69.6
+DeiT-III-B [54] 86M 17.6G 77.6 84.8 56.6 70.3
+ViL-B 89M 18.6G 76.6 84.7 59.1 70.9
+Table 3: Transfer classification accuracies on the VTAB-1K [64] benchmark using ImageNet-1K pre-trained models.
+VTAB-1K consists of 19 datasets split into 7 natural, 4 specialized and 8 structured datasets.
+We show averages per category and the average accuracy over all 19 datasets (Appendix Table 10 lists all individual accuracies).
+ViL shows strong transfer classification performance, outperforming heavily optimized ViT protocols and Vim on the full VTAB-1K benchmark.
+Notably, ViL performs exceptionally well on the structured category.
+We tune the learning rate for each model and dataset on the validation set and report the average performance on the test set over 5 seeds.
+Appendix Table 13 lists further hyperparameters.
+3.4 Block Design
 We study different ways to design ViL blocks in Table 4.
@@
 Figure 3 visualizes the different traversal paths.
-Figure 2: Visualization of traversal paths. Bi-directional uses the first two traversal paths. Quad-directional uses all of the above. Squares represent individual patch tokens.
 Due to the increased cost of bi-directional and quad-directional blocks, this study was conducted in a heavily reduced setting.
@@
 Since quad-directional designs are compute intensive, this study was conducted by training a ViL-T on an ImageNet-1K subset of 100 classes using resolution 128.
+“Param.
+Sharing” shares the parameters of the two (bidirectional) or four (quaddirectional) mLSTM layers in each block.
+Figure 3: Visualization of traversal paths. Bi-directional uses the first two traversal paths. Quad-directional uses all of the above. Squares represent individual patch tokens.
 3.5 Classification Design
@@
 [AVG] averages all patch tokens, “Middle Patch” uses the middle patch token, “Middle [CLS]” uses a [CLS] token in the middle of the sequence, “Bilateral [AVG]” uses the average of the first and last patch token.
-We find ViL to be relatively robust the classification design where all performances are within 0.6%.
-We choose “Bilateral [AVG]” over “Middle [CLS]”, as ImageNet-1K is known to have a center bias, i.e., objects are commonly located in the middle of the picture.
-By using “Bilateral [AVG]” we avoid exploiting this bias, which makes our model more general.
-Arch ACC
-[AVG] 75.4
-Middle Patch 75.9
-Middle [CLS] 75.3
-Bilateral [AVG] 75.9
-Bilateral Concat 76.4
-Table 3: How to pool the sequence of patch tokens into a single token for classification?
+We find ViL to be relatively robust the classification design where all performances that use a single token are within 0.6%.
+We only add a [CLS] token if it is used for classification, hence the best design (“Bilateral Concat”) does not use a [CLS] token.
+Design Dim ACC
+[AVG] 192 75.4
+Middle Patch 192 75.9
+Middle [CLS] 192 75.3
+Bilateral [AVG] 192 75.9
+Bilateral Concat 384 76.4
+Table 5: How to pool the sequence of tokens for classification?
 ViL is relatively robust to the classification design.
@@
 The good performance of middle patch could be caused by the bias in ImageNet-1K that objects are typically in the center of the image.
-We therefore choose “Bilateral [AVG]” to avoid exploiting this bias.
-“Bilateral Concat” concatenates the first and last patch token which introduces additional parameters in the classification head.
-To save compute, we conduct this study by training a ViL-T on ImageNet-1K for 400 epochs in 192x192 resolution followed by fine-tuning for 20 epochs on 224x224 resolution.
-To remain comparable to previous architectures that use a single token as input to the classification head, we average the first and the last patch.
-To achieve best performances, we suggest to concatenate the two tokens (“Bilateral Concat”) instead of averaging them.
-This is similar to common practices in self-supervised vision transformers, such as DINOv2 [28], which are trained with two objectives that are attached at the [CLS] and [AVG] token respectively and therefore benefit from concatenating the representation of [CLS] and [AVG] token.
-This directions has also been explored for vision SSM models [40] where multiple [CLS] tokens are spread out over the sequence and then used as input to a classifier.
-Similar directions could also improve the performance of ViL.
+We therefore choose a bilateral design to avoid exploiting this bias.
+Additionally, concatenating the first and last token (“Bilateral Concat”) instead of averaging improves performance and is used as final design for ViL.
+We conduct this study by training a ViL-T on ImageNet-1K for 400 epochs.
+3.6 Vision Adjustments of xLSTM
+The xLSTM [5] architecture was designed for language modeling and therefore includes some design choices that are specific to language modeling.
+One of these design choices is a causal 1D convolution (with kernel size 4) to encode local context into the query and key input of the mLSTM layer.
+As images are neither causal nor 1D structures, we replace the causal 1D convolution with a 2D convolution (with kernel size 3).
+When training a ViL-T for 400 epochs, this improves ImageNet-1K accuracy from 76.1% to 76.7%.
+Another language specific design choice is the absence of bias parameters in projection and layernorm [4] layers which improves training stability for larger models [12, 16].
+As vision models are typically much smaller than language models, ViTs commonly use all bias parameters.
+When adding bias parameters, a ViL-T trained for 400 epochs improves from 76.7% to 77.1% on ImageNet-1K.
 4 Conclusion
@@
 Odd blocks process image patches row-wise from top left to bottom right and even blocks go from bottom right to top left.
-Our new architecture outperforms SSM-based vision architectures and also optimized ViT models on ImageNet-1K classification.
-Remarkably, ViL is able to outperform ViT training pipelines, that resulted from years of hyperparameter tuning and transformer improvements, in fair comparisons.
+Our new architecture outperforms SSM-based vision architectures and also optimized ViT models on ImageNet-1K classification, VTAB-1K transfer classification and ADE20K semantic segmentation.
+Remarkably, ViL is able to outperform ViT training pipelines, which are the result of years of hyperparameter tuning and transformer improvements.
 In the future, we see potential in applying ViL when high-resolution images are needed for optimal performance, such as semantic segmentation or medical imaging.
-In these settings, transformers suffer from high computational costs due to the quadratic complexity of self-attention, which ViL does not have due to its linear complexity.
+In these settings, transformers suffer from high computational costs due to the quadratic complexity of self-attention, where ViL can use a chunked form to trade-off between its parallel form (quadratic complexity) and recurrent form (linear complexity).
 Additionally, improving pre-training schemes (e.g.
-via self-supervised learning), exploring better hyperparameter settings or transferring techniques from transformers (e.g.
-LayerScale [35]) are promising directions of ViL.
+via self-supervised learning), exploring better hyperparameter settings or investigating hierarchical architectures are promising future directions that could improve ViL even further.
 Acknowledgments
@@
 Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.
-Appendix A Extended Results
-A.1 Runtime Comparison of ViL vs Vim
+Appendix A Changelog
+A.1 Version 1
+In the first version of this paper, we trained models with a causal 1D convolution in the mLSTM layer (instead of a 2D convolution), “Bilateral [AVG]” pooling (instead of “Bilateral Concat”) and a constant resolution of 224 (instead of pre-training at 192 followed by a short fine-tuning on resolution 224).
+This version of the architecture is also used to conduct the studies in Table 4 and Table 5.
+For completeness, we report the accuracies of these models in Table 6.
+In the first version, we trained slightly deeper models (ViL-T+ with 29 blocks and ViL-S+ with 26 blocks) to match the parameter count of Vim-T and Vim-S.
+However, as the models with a regular depth of 24 were already better than Vim with less parameters, we omit them in future versions.
+Method #Params ACC
+ViL-T 6M 77.3
+ViL-T+ 7M 78.1
+ViL-S 23M 80.7
+ViL-S+ 26M 80.9
+ViL-B 88M 81.6
+Model #Params ACC
+Long-sequence fine-tuning
+ViL-T+ 7M 80.0
+ViL-S+ 26M 82.2
+ViL-B 88M 82.7
+Table 6: ImageNet-1K accuracies of models trained in the first version of this paper.
+A.2 Version 2
+The current version of the paper improves performance by (i) replacing the causal 1D convolution with a 2D convolution (with kernel size 3 and padding 1) (ii) using “Bilateral Concat” instead of “Bilateral [AVG]” pooling (iii) using bias parameters in projection and LayerNorm [4] layers (iv) pre-training models on 192x192 followed by a short fine-tuning on 224x224 resolution.
+Appendix B Extended Results
+B.1 Runtime Comparison of ViL vs Vim
 We compare the time it takes to train ViL and Vim [67] for a single ImageNet-1K epoch in Table 7.
 We follow the scaling procedure of ViTs, using 192 (T), 384 (S), 768 (B), 1024 (L) as latent dimension where the large scale additionally doubles the number of blocks.
-Optimization (T)iny (S)mall (B)ase (L)arge
+Model Optimization (T)iny (S)mall (B)ase (L)arge
 Vim [67] custom CUDA kernel 7.3h 14.0h 28.2h 76.4h
@@
 Runtimes denote the training time for 10 ImageNet-1K epochs and are extrapolated from short benchmark runs on a single A100-80GB-PCIe using float16 precision and 224x224 images.
-A.2 Impact of Longer Training
-We investigate training for a longer duration on ViL-T+ and train it for 400 and 800 epochs following the hyperparameters in Table 5.
-After 400 epochs, ViL-T+ achieves 77.2% accuracy and after doubling the training duration the model achieves 78.1% accuracy.
-Our implementation of DeiT-III-T achieves 75.6% after 400 epochs and 76.2% with 800 epochs using the hyperparameters in Table 6.
-Appendix B Implementation Details
-B.1 Hardware
+B.2 Impact of Longer Training
+We investigate the impact of training for a longer duration in Table 8.
+We also attempted to train larger models for longer but did not observe similar performance gains.
+One could probably change that by tuning hyperparameters for an 800 epoch training schedule, which we omit due to the large computational costs associated with training larger models.
+Model Epochs IN-1K ACC VTAB-1K ADE20K mIoU
+DeiT-III-T 400 75.6 67.0 39.1
+DeiT-III-T 800 76.2 67.1 39.8
+ViL-T 400 77.2 67.8 40.9
+ViL-T 800 78.3 68.3 41.2
+Table 8: Performance comparison of tiny models when training for 400 and 800 epochs. ADE20K mIoU uses single-scale evaluation. All settings follow the ones used in the main paper.
+B.3 Long-sequence fine-tuning
+Table 9 shows results for the long-sequence fine-tuning setting of Vim [67].
+Method #Params ACC
+DeiT-III-T (reimpl.) 6M 78.3
+Vim-T [67] 7M 78.3
+ViL-T 6M 80.1
+DeiT-S [51] 22M 81.4
+DeiT-III-S [54] 22M 83.2
+Vim-S [67] 26M 81.6
+ViL-S 23M 82.7
+DeiT-B [51] 86M 83.2
+DeiT-III-B [54] 86M 84.5
+ViL-B 89M 83.3
+Table 9: Fine-tuning ImageNet-1K pre-trained models by using a 50% overlap when creating patches.
+This setting increases the number of patch tokens from 196 to 729.
+ViL easily outperforms Vim but falls short of ViTs on larger models.
+Note that ViL FLOPS are calculated using the parallel form of the mLSTM, which scales quadratically with the sequence length.
+Using the chunkwise form instead of the parallel form can reduce FLOPS for longer sequences (see Section C.2).
+B.4 VTAB-1K Individual Dataset Results
+Table 10 presents accuracies for each individual dataset of the VTAB-1K [64] benchmark.
+Natural Specialized Structured
+Cifar100
+Caltech101
+DTD
+Flower102
+Pets
+SVHN
+Sun397
+Camelyon
+EuroSAT
+Resisc45
+Retinopathy
+Clevr-Count
+Clevr-Dist
+DMLab
+KITTI-Dist
+dSpr-Loc
+dSpr-Ori
+sNORB-Azim
+sNORB-Ele
+DeiT-T [51] 47.7 86.4 63.7 85.6 87.0 78.4 35.3 83.0 93.4 80.9 70.7 71.7 60.3 43.1 78.5 67.9 41.6 30.6 32.7
+DeiT-III-T (reimpl.) 52.3 90.1 62.7 88.8 87.5 83.7 37.9 83.2 93.1 81.1 72.9 76.6 60.8 44.9 79.1 67.5 48.1 31.0 33.3
+Vim-T [67] 46.7 86.3 60.7 84.0 88.8 76.1 33.7 82.2 92.9 75.2 72.6 59.8 49.9 39.3 78.2 51.2 43.9 26.9 27.2
+ViL-T 54.2 90.2 67.4 90.7 89.9 81.6 41.1 83.4 94.2 82.7 73.1 80.7 61.8 49.4 81.3 57.8 51.8 31.4 34.8
+DeiT-S [51] 57.0 88.9 68.2 90.9 90.8 75.4 42.1 83.3 94.0 83.8 74.0 74.6 58.3 45.6 78.2 61.9 47.9 27.1 31.9
+DeiT-III-S [54] 58.8 88.6 67.5 90.9 91.7 84.4 43.3 84.4 92.6 82.5 73.5 76.5 57.9 46.2 78.9 58.3 49.7 23.7 27.5
+Vim-S [67] 53.0 87.2 64.1 86.8 90.3 65.8 39.7 82.4 93.4 78.0 73.1 63.1 53.2 42.3 78.2 54.1 47.6 27.1 29.3
+ViL-S 61.4 89.6 69.2 92.8 91.7 78.7 43.8 85.5 93.9 84.4 73.5 84.0 63.4 51.3 83.3 61.0 55.4 32.4 35.5
+DeiT-B [51] 61.8 89.8 67.5 93.7 92.6 84.4 45.6 85.3 95.1 86.3 74.2 77.7 59.9 47.2 81.7 61.7 51.4 30.0 36.2
+DeiT-III-B [54] 62.9 89.6 69.6 93.7 93.2 87.0 47.1 85.8 94.1 85.6 73.7 80.5 61.4 48.4 80.9 64.4 55.1 30.2 31.8
+ViL-B 64.3 90.0 71.1 93.4 91.4 79.6 46.6 84.7 94.3 85.4 74.4 83.7 62.1 52.7 81.0 63.1 57.6 32.6 39.9
+Table 10: Results on all datasets of the VTAB-1K [64] benchmark.
+B.5 Robustness
+Table 11 presents robustness and OOD evaluations of ImageNet-1K pre-trained classifiers.
+Model IN-C ($\downarrow$) IN-A ($\uparrow$) IN-R ($\uparrow$) Sketch ($\uparrow$) Validation ($\uparrow$)
+DeiT-T [51] 69.7 7.6 32.7 19.9 72.2
+DeiT-III-T 65.0 11.7 39.4 27.4 76.2
+Vim-T [67] 61.8 9.6 38.8 26.9 76.1
+ViL-T 59.6 15.2 42.2 30.0 78.3
+DeiT-S [51] 54.4 19.6 41.9 29.1 79.8
+DeiT-III-S [54] 50.1 23.2 46.6 35.4 81.4
+Vim-S [67] 51.5 19.7 44.8 32.5 80.5
+ViL-S 50.6 23.8 47.9 35.2 81.5
+DeiT-B [51] 48.6 27.9 44.6 32.0 81.8
+DeiT-III-B [54] 42.7 36.5 54.1 41.1 83.8
+ViL-B 45.3 30.9 51.9 39.0 82.4
+Table 11: Robustness and OOD evaluations on ImageNet-C(orruption) [29], ImageNet-A(dversarial) [31], ImageNet-R(endition) [30] and ImageNet-Sketch [59]..
+For ImageNet-C, we report the mean corruption error [29] with AlexNet [39] as baseline.
+Appendix C Implementation Details
+C.1 Hardware
 We train models on a mix of custom hardware servers (mainly A100 and A40 GPUs) and public research clusters equiped with either 8xA100 or 4xA100 nodes.
-We estimate the total number of A100 GPU-hours used for this project to be 25K hours. This estimate includes everything from initial exploration, method development, analysis and evaluations.
-B.2 FLOPS Calculation
-We calculate the FLOPS from Table 2 with the fvcore library [^1] https://github.com/facebookresearch/fvcore.
-As fvcore does not support all operations of ViL, the FLOPS are not 100% accurate, but nevertheless serve as a good reverence point to compare the relative computational effort between block designs.
-B.3 ViL Hyperparameters
+We estimate the total number of A100 GPU-hours used for this project to be 38K hours. This estimate includes initial exploration, method development, analysis and evaluations.
+C.2 FLOPS Calculation
+We calculate the FLOPS with the fvcore library [^1] https://github.com/facebookresearch/fvcore.
+For ViL, we report the FLOPS of the mLSTM parallel form.
+Given sequence length $L$ and hidden dimension $D$, it has complexity $\mathcal{O}(L\frac{L}{2}D)$ where the $\frac{L}{2}$ stems from the causal masking.
+We report FLOPS for a complexity of $\mathcal{O}(L\frac{L}{2}D)$ even though our current implementation of the mLSTM has complexity $\mathcal{O}(L^{2}D)$ (i.e., entries of the $QK$ matrix that are later masked are still calculated) which we justify by the fact that FlashAttention-2 [14] is approximately 1.7x faster with a causal mask than without.
+Therefore, an optimized hardware implementation of the mLSTM could also omit the calculation of values that are not needed.
+As Vim [67] does not report FLOPS and their model makes use of CUDA kernels (which are not counted as FLOPS by fvcore), we replace all calls to CUDA kernels with their reference PyTorch implementation and count the FLOPS with fvcore.
+For the total pre-training compute in Figure 2, we consider an efficient implementation of stochastic depth [32] which omits the calculation of a dropped block instead of masking it.
+Therefore, we change the implementation of ViT [18] to use our efficient stochastic depth implementation.
+Vim does not use stochastic depth for training as their models are relatively small.
+C.3 ViL Hyperparameters
 Table 12 shows detailed hyperparameters used to train ViL models.
-In the “long-sequence fine-tuning” setting, we follow Vim [44] and fine-tune for 30 epochs with a learning rate of 1e-5, a weight decay of 1e-8 using an overlap of 50% (8 pixels) between consecutive patches.
-This increases the sequence length from 196 to 729.
-Other hyperparameters follow Table 5.
 Parameter Value
-Epochs 800 (T), 400 (S/B)
-Batch size 1024 (S), 2048 (T/B)
+Epochs 800 (T), 400 (S/B) $\rightarrow$ 20 (T, S), 5 (B)
+Batch size 2048 $\rightarrow$ 1024
 Model
@@
    Latent dimension 192 (T), 384 (S), 768 (B)
-   Depth 24 (T, S, B), 29 (T+), 26 (S+)
+   Depth 24
+   Pooling Bilateral Concat
 Stochastic depth [32]
@@
 Optimizer AdamW [42, 36]
-   Base Learning rate 1e-3
+   Base Learning rate 1e-3 $\rightarrow$ 1e-5
    Linear LR Scaling Divisor [22] 1024
@@
    Gradient Norm Clip 1.0
-Learning rate schedule linear warmup $\rightarrow$ cosine decay
-   Warmup epochs 5
+Precision mixed bfloat16 [35]
+   Backend torch.autocast
+Learning rate schedule cosine decay
+   Warmup schedule linear
+   Warmup epochs 5 $\rightarrow$ 5 (T, S), 1 (B)
    End LR 1e-6
@@
 Train Data Augmentation
-   RandomResizedCrop 224
+   RandomResizedCrop 192 $\rightarrow$ 224
     Scale [0.08, 1.0]
@@
 Test Data Augmentation
-   Resize 224
+   Resize 192 $\rightarrow$ 224
     Interpolation bicubic
-   CenterCrop 224
+   CenterCrop 192 $\rightarrow$ 224
    Normalize ImageNet-1K statistics
 Table 12: Hyperparameters for training ViL on ImageNet-1K, inspired by DeiT-III [54].
-B.4 DeiT-III Reimplementation Hyperparameters
+We follow the best setting from DeiT-III [54] and pre-train on 192 resolution followed by a short fine-tuning on 224 resolution (indicated by $\rightarrow$).
+C.4 Fine-tuning on VTAB-1K
+For fine-tuning models on VTAB-1K we provide the hyperparameters in Table 13.
+We search for the best learning rate for each dataset by fine-tuning the model 25 times (5 learning rates with 5 seeds each) on the 800 training samples and evaluating them on the 200 validation samples.
+With the best learning rate, we then train each model 5 times on concatenation of training and validation split, evaluate on the test split and report the average accuracy.
+Parameter Value
+Epochs 50
+Batch size 64
+Seeds 5
+Optimizer AdamW [42, 36]
+   Learning rate [1e-3, 7.5e-4, 5.0e-4, 2.5e-4, 1.0e-4]
+   Layer-wise lr decay [13] 0.65*
+   Weight decay 0.05
+   Momentum $\beta_{1}=0.9,\beta_{2}=0.999$
+Learning rate schedule linear warmup $\rightarrow$ cosine decay
+   Warmup epochs 5
+Precision mixed bfloat16 [35]
+   Backend torch.autocast
+Data Augmentation
+   Resize
+    interpolation bicubic
+    size 224x224
+   Normalize ImageNet-1K statistics
+Table 13: Hyperparameters for fine-tuning on VTAB-1K.
+*For Vim and ViL we group two consecutive blocks for the layer-wise lr decay similar to how ViT considers a pair of attention and MLP block as a single “layer” for the decay.
+C.5 ADE20K Semantic Segmentation Fine-tuning
+We fine-tune models on ADE20K [66] using an UperNet [61] head.
+We follow common practices and fine-tune on 512x512 resolution, where we interpolate the absolute positional embedding from 224x224 to 512x512.
+For ViTs, we add relative position biases to the attention layers (initialized to 0) [27].
+Table 14 lists detailed hyperparameters.
+Parameter Value
+Updates 160K
+Batch size 16
+UperNet
+   Auxiliary
+    Weight 0.4
+    Input Block 8*
+    Dimension 192 (T), 384 (S, B)
+   Decoder
+    Weight 1.0
+    Input Blocks [4, 6, 8, 12]*
+    Dimension 192 (T), 384 (S, B)
+Stochastic depth [32]
+   Peak rate 0 (T), 0.05 (S), 0.1 (B)
+   Layer-wise Decay
+Optimizer AdamW [42, 36]
+   Learning rate 5e-4
+   Linear LR Scaling Divisor [22] 16
+   Layer-wise lr decay [13] 0.65*
+   Weight decay 0.05
+   Momentum $\beta_{1}=0.9,\beta_{2}=0.999$
+Learning rate schedule linear warmup $\rightarrow$ cosine decay
+   Warmup updates 1500
+Precision mixed float16
+   Backend torch.autocast
+Train Data Augmentation
+   RandomResize
+    interpolation bicubic
+   RandomCrop
+    size 512x512
+   RandomHorizontalFlip
+   ColorJitter 0.5
+    brightness 0.5
+    contrast 0.5
+    saturation 0.5
+    hue 0.25
+   Normalize ImageNet-1K statistics
+Evaluation
+   Stride 341
+   Multi-scale
+    scale factors [0.75, 1.0, 1.25, 1.5, 1.75]
+    flip [True, False]
+Table 14: Hyperparameters for fine-tuning on VTAB-1K. *For ViL we group two consecutive blocks into one similar to how a ViT block consists of a pair of attention and MLP block.
+C.6 DeiT-III Reimplementation Hyperparameters
 Table 12 shows detailed hyperparameters used to train DeiT-III-T (reimpl.) from Table 1.
@@
 Epochs 800 $\rightarrow$ 20
-Batch size 2048
+Batch size 2048 $\rightarrow$ 1024
+Model
+   Patch size 16x16
+   Latent dimension 192
+   Depth 12
+   Pooling [CLS]
 Stochastic depth [32]
@@
 Optimizer AdamW [42, 36]
-   Base Learning rate 1e-3
+   Base Learning rate 1e-3 $\rightarrow$ 1e-5
    Linear LR Scaling Divisor [22] 1024
@@
    Gradient Norm Clip
-Learning rate schedule linear warmup $\rightarrow$ cosine decay
+Precision mixed bfloat16 [35]
+   Backend torch.autocast
+Learning rate schedule cosine decay
+   Warmup schedule linear
    Warmup epochs 5

### v3

@@
 Recently, the Long Short-Term Memory (LSTM) has been extended to a scalable and performant architecture – the xLSTM – which overcomes long-standing LSTM limitations via exponential gating and parallelizable matrix memory structure.
-In this report, we introduce Vision-LSTM (ViL), an adaption of the xLSTM building blocks to computer vision.
+In this paper, we introduce Vision-LSTM (ViL), an adaption of the xLSTM building blocks to computer vision.
 ViL comprises a stack of xLSTM blocks where odd blocks process the sequence of patch tokens from top to bottom while even blocks go from bottom to top.
+ViL achieves strong performances on classification, transfer learning and segmentation tasks as well as a beneficial pre-training cost-to-performance trade-off.
 Experiments show that ViL holds promise to be further deployed as new generic backbone for computer vision architectures.
 Project page: https://nx-ai.github.io/vision-lstm/
+Figure 1: The efficient and scalable design of Vision-LSTM shows strong performances, uses less FLOPS than Transformer/Mamba counterparts and scales linear to higher resolutions.
+Performance is averaged over ImageNet accuracy, ADE20K mIoU and VTAB-1K accuracy.
+footnotetext: Published as a conference paper at ICLR 2025
 1 Introduction
@@
 This simple alternating design allows ViL to efficiently process non-sequential inputs, such as images, without introducing additional computations.
-Similar to vision adaptions of SSMs [40, 67, 58], ViL can exhibit linear computational and memory complexity w.r.t.
-sequence length which makes it appealing for tasks that benefit from high-resolution images such as medical imaging [9, 26, 55, 62], segmentation [37, 10], or physics simulations [6, 43, 7, 2].
-In contrast, ViT’s computational complexity scales quadratically due to the self-attention mechanism, rendering them costly to apply to high-resolution tasks.
-2 Method
-Vision-LSTM (ViL) is a generic backbone for computer vision tasks, which is residually built from xLSTM blocks, as visualized in Figure 1.
-Following ViT [18], ViL first splits an image into non-overlapping patches via a shared linear projection, then adds learnable positional embeddings to each patch token.
-At the core of ViL are alternating mLSTM blocks, which are fully parallelizable and equipped with a matrix memory combined with a covariance update rule.
-Odd mLSTM blocks process patch tokens from top left to bottom right while even blocks go from bottom right to top left.
 Figure 2: Schematic overview of Vision-LSTM (ViL).
@@
 For classification, ViL uses the concatenation of the first and the last patch as input to a linear classification head.
+ViL is an isotropic architecture, i.e., all blocks have the same input and output dimension and no downsampling layers are used except the initial patch embedding.
+Projection layers process each patch individually and the mLSTM exchanges information between patches.
+Similar to vision adaptions of SSMs [49, 82, 73], ViL can exhibit linear computational and memory complexity w.r.t.
+sequence length which makes it appealing for tasks that benefit from high-resolution images such as medical imaging [10, 32, 69, 77], segmentation [44, 13], or physics simulations [6, 52, 7, 2].
+In contrast, ViT’s computational complexity scales quadratically due to the self-attention mechanism, rendering them costly to apply to high-resolution tasks.
+Our contributions summarize as follows:
+  • We introduce Vision-LSTM (ViL), an adaption of the mLSTM to computer vision tasks that can serve as a generic vision backbone with linear complexity.
+  • We show modeling capacity and generalization in the common vision benchmark of pre-training models on ImageNet-1K, followed by fine-tuning on transfer classification and semantic segmentation tasks.
+  • We ablate various architectural design choices to evaluate their impact on performance and provide insights into the model design.
+  • We discuss potential future directions and current limitations that, once addressed, will improve ViL even further.
+2 Method
+Vision-LSTM (ViL) introduces xLSTM [5] to computer vision, similar to other vision adaptions of sequence modeling architectures, e.g., Vision Transformers [22], Vision Mamba [82], or Vision RWKV [23].
+2.1 Preliminaries
+[…]
+The key innovations of the mLSTM [5] are the enhanced storage capacity compared to the classical LSTM [39] by using a matrix memory cell ${\bm{C}}\in\mathbb{R}^{d\times d}$ instead of a scalar memory cell $c\in\mathbb{R}$ and introducing exponential gates (instead of sigmoid gates) to the input and forget gates, where $d$ is the hidden dimension within the mLSTM block (typically $d=2D$).
+Intuitively, the mLSTM is a more expressive and faster version of the classical LSTM that can be efficiently parallelized on modern hardware.
+In ViL, the mLSTM is used to process dependencies between patches, similar to how the attention exchanges information between patches in a ViT.
+The mLSTM is embedded into a gated MLP architecture, as shown on the right of Figure 2, where the weight matrices of the MLP process each patch individually and the mLSTM exchanges information between patches.
+For completeness, we outline the forward pass of the mLSTM in the following paragraphs.
+The mLSTM [5] is a recurrent neural network, which maps a state $({\bm{h}}_{t-1},{\bm{C}}_{t-1},{\bm{n}}_{t-1})$ to a successor state $({\bm{h}}_{t},{\bm{C}}_{t},{\bm{n}}_{t})$ given input ${\bm{x}}_{t-1}$.
+Thereby, ${\bm{h}}_{t}\in\mathbb{R}^{d}$ denotes the hidden state, ${\bm{C}}_{t}\in\mathbb{R}^{d\times d}$ is the cell state and ${\bm{n}}_{t}\in\mathbb{R}^{d}$ corresponds to a normalizer state.
+The full forward pass of the mLSTM is as follows [5]:
+${\bm{C}}_{t}$ $=\ f_{t}\ {\bm{C}}_{t-1}\ +\ i_{t}\ {\bm{v}}_{t}\ {\bm{k}}_{t}^{\top}$ cell state (1)
+${\bm{n}}_{t}$ $=\ f_{t}\ {\bm{n}}_{t-1}\ +\ i_{t}\ {\bm{k}}_{t}$ normalizer state (2)
+${\bm{h}}_{t}$ $=\ {\bm{o}}_{t}\ \odot\ \tilde{{\bm{h}}}_{t}$ $\tilde{{\bm{h}}}_{t}$ $=\ {\bm{C}}_{t}{\bm{q}}_{t}\ /\ \max\left\{|{{\bm{n}}_{t}^{\top}{\bm{q}}_{t}}|,1\right\}$ hidden state (3)
+${\bm{q}}_{t}$ $=\ {\bm{W}}_{q}\ {\bm{x}}_{t}\ +\ {\bm{b}}_{q}$ query input (4)
+${\bm{k}}_{t}$ $=\ \frac{1}{\sqrt{d}}{\bm{W}}_{k}\ {\bm{x}}_{t}\ +\ {\bm{b}}_{k}$ key input (5)
+${\bm{v}}_{t}$ $=\ {\bm{W}}_{v}\ {\bm{x}}_{t}\ +\ {\bm{b}}_{v}$ value input (6)
+$i_{t}$ $=\ \exp\!\big{(}\tilde{i}_{t}\big{)}$ $\tilde{i}_{t}$ $=\ {\bm{w}}^{\top}_{i}\ {\bm{x}}_{t}\ +\ b_{i}$ input gate (7)
+$f_{t}$ $=\ \exp\!\big{(}\tilde{f}_{t}\big{)}$ $\tilde{f}_{t}$ $=\ {\bm{w}}^{\top}_{f}\ {\bm{x}}_{t}\ +\ b_{f}$ forget gate (8)
+${\bm{o}}_{t}$ $=\ \sigma\big{(}\tilde{fo}_{t}\big{)}$ $\tilde{{\bm{o}}}_{t}$ $=\ {\bm{W}}_{{\bm{o}}}\ {\bm{x}}_{t}\ +\ {\bm{b}}_{{\bm{o}}}$ output gate (9)
+As exponential activation functions can lead to large activations, the input and forget gates are stabilized with an additional state $m_{t}$:
+$m_{t}$ $=\max\!\Big{(}\log(f_{t})+{\bm{m}}_{t-1},\log(f_{t})\Big{)}$ stabilizer state (10)
+$i^{\prime}_{t}$ $=\exp\!\Big{(}\log(i_{t})-m_{t}\Big{)}=\exp\Big{(}\tilde{i}-m_{t}\Big{)}$ stabilized input gate (11)
+$f^{\prime}_{t}$ $=\exp\!\Big{(}\log(f_{t})+m_{t-1}-m_{t}\Big{)}$ stabilized forget gate (12)
+As the mLSTM has no memory mixing, i.e, interactions between hidden states from one timestep to the next, it can be fully parallelized for fast computation on modern hardware.
+For a detailed discussion and theory of the cell state update, further details to the mLSTM we refer to the original work [5].
+2.2 Vision-LSTM (ViL)
+Vision-LSTM (ViL) is a generic backbone for computer vision tasks, which is residually built from mLSTM blocks, as visualized in Figure 2.
+Following ViT [22], ViL first splits an image into non-overlapping patches via a shared linear projection, then adds learnable positional embeddings to each patch token.
+At the core of ViL are alternating mLSTM blocks, which are fully parallelizable and equipped with a matrix memory combined with a covariance update rule.
+Odd mLSTM blocks process patch tokens from top left to bottom right while even blocks go from bottom right to top left.
+Formally, the forward pass of a pair of ViL blocks is:
+${\bm{Y}}^{\prime}$ $={\bm{X}}+\text{Block}_{\theta}({\bm{X}})$ (13)
+${\bm{Y}}$ $={\bm{Y}}^{\prime}+\text{Flip}(\text{Block}_{\phi}(\text{Flip}({\bm{Y}}^{\prime})))$ (14)
+Where “Flip” reverses the sequence and “$\text{Block}_{\theta}$” and “$\text{Block}_{\phi}$” corresponds to mLSTM blocks with parameters $\theta$ and $\phi$ (shown in Figure 2, right).
+A key motivation of ViL is that the autoregressive mLSTM can operate in a recurrent, parallel or chunkwise mode, each with distinct FLOPS and runtime characteristics.
+Given a sequence length $T$ and hidden dimension $d$, the complexity of the recurrent mode is $\mathcal{O}(Td^{2})$ and needs to be processed sequentially, whereas the parallel mode has complexity $\mathcal{O}(T^{2}d)$ and is fully parallelizable.
+The chunkwise mode combines the advantages of the other modes by introducing a chunksize $S$ where the parallel mode is used within chunks and the recurrent mode between chunks.
+This allows high parallelization, minimal operations and linear scaling with $T$.
+Complexity wise, the chunkwise mode has $\mathcal{O}(\frac{T}{S}S^{2}d+\frac{T}{S}d^{2})$ or $\mathcal{O}(TSd+\frac{T}{S}d^{2})$ where $\frac{T}{S}$ corresponds to the number of chunks.
 3 Experiments
+Figure 3: Performance overview of ImageNet-1K pre-trained models in relation to pre-training compute.
+ViL shows strong performances across classification (ImageNet-1K), semantic segmentation (ADE20K) and transfer classification (VTAB-1K) tasks.
 We pre-train models on ImageNet-1K [19], which contains 1.3M training images and 50K validation images where each image belongs to one of 1000 classes.
@@
 Detailed hyperparameters can be found in Appendix Table 10.
-As ViL is an isotropic architecture (like ViT [18]), we compare it against other isotropic architectures and leave exploration of hierarchical ViL architectures to future work.
-Figure 2: Performance overview of ImageNet-1K pre-trained models in relation to pre-training compute. ViL shows strong performances across classification and semantic segmentation tasks.
-As ViTs are well established in the vision community, they underwent multiple optimization cycles over the years [18, 51, 53, 52, 54, 27].
-Since this work is the first to apply xLSTM to computer vision, we do not expected to outperform years of hyperparameter tuning of ViTs in all cases.
-Even so, Figure 2 shows an overview of performance metrics in relation to total pre-training compute where ViL performs favorably against heavily optimized transformer protocols (DeiT, DeiT-III) and Vision Mamba (Vim) [67] on ImageNet-1K [17] classification, ADE20K [66] semantic segmentation and VTAB-1K [64] classification.
+We then transfer the pre-trained models to serveral benchmark tasks: ImageNet-1K classification on the validation set, ADE20K [81] semantic segmentation and VTAB-1K [79] classification.
+These benchmarks evaluate global image understanding (ImageNet-1K), semantic local and global understanding (ADE20K) and few-shot generalization to a diverse set of 19 VTAB-1K classification datasets, which include natural images, specialized imagery (medical and satellite) and structured tasks (camera angle prediction, depth estimation, object counting, …).
+Figure 3 shows an overview of performance metrics in relation to total pre-training compute where ViL performs favorably against heavily optimized transformer protocols (DeiT, DeiT-III) and Vision Mamba (Vim).
 Detailed results are presented in the following sections.
-3.1 ImageNet-1K Pre-training
-Table 1 relates parameter counts and FLOPS to classification accuracy after pre-training on ImageNet-1K.
+As ViTs are well established in the vision community, they underwent multiple optimization cycles over the years [22, 64, 66, 65, 67].
+Therefore, a vast part of the hyperparameter space for pre-training ViTs has been explored.
+Since this work is the first to apply xLSTM to computer vision, considerably less effort has been put into hyperparameter tuning and architecture optimization, suggesting that future work could improve ViL even further.
+3.1 ImageNet-1K Classification
+Table 1 relates parameter counts and FLOPS to validation accuracy after pre-training on ImageNet-1K.
 ViL outperforms heavily optimized ViT protocols and other backbones on the tiny and small scale.
-While ViL does not outperform all other models on the base scale, evaluations on downstream tasks (as shown later in Table 3 and Table 2) show that ViL-B still learns strong features.
-Additionally, training a ViL-B for 400 epochs takes roughly 600 A100 GPU-hours or 19 hours on 32 A100 GPUs.
-Therefore, the hyperparameter configuration is most likely not optimal as we were not able to extensively optimize hyperparameters for this scale.
-Note that training ViL is still quite expensive to train as there does not exist an optimized hardware implementation yet, which does exist for other models like ViT [15, 14] or Vim [23, 67].
-Nevertheless, the efficient alternating block design of ViL makes it already faster than Vim (up to 69% as shown in Appendix B.1) while only using torch.compile, a generic speed optimization of the PyTorch [45] framework.
+While ViL does not outperform all other models on the base scale, evaluations on downstream tasks (as shown later in Table 2 and Table 3) show that ViL-B still learns strong features, particularly for semantic segmentation and structured tasks.
 Model Epochs #Params FLOPS IN-1K
@@
 Vim-T [82] 300 7M 1.5G 76.1
-Mamba®-T [58] 280+20 9M - 77.4
-ViL-T 800+20 6M 1.5G 78.3
+Mamba®-T [73] 280+20 9M 1.6G 77.4
+ViL-T 800+20 6M 1.3G 78.3
 DeiT-S [64] 300 22M 4.6G 79.8
@@
 DeiT-III-S [67] 800+20 22M 4.6G 81.4
-ConvNeXt-S [41] (iso.) 300 22M 4.3G 79.7
+ConvNeXt-S (iso.[50] 300 22M 4.3G 79.7
 VRWKV-S [23] 300 24M 4.6G 80.1
 Vim-S [82] 300 26M 5.3G 80.5
-Mamba®-S [58] 280+20 28M - 81.1
-ViL-S 400+20 23M 5.1G 81.5
+Mamba®-S [73] 280+20 28M 5.5G 81.1
+ViL-S 400+20 23M 4.7G 81.5
 DeiT-B [64] 300 86M 17.6G 81.8
@@
 DeiT-III-B [67] 800+20 86M 17.6G 83.7
-ConvNeXt-B [41] (iso.) 300 87M 16.9G 82.0
+ConvNeXt-B (iso.[50] 300 87M 16.9G 82.0
 VRWKV-B [23] 300 94M 18.2G 82.0
-Mamba®-B [58] 280+20 99M - 82.9
-ViL-B 400+5 89M 18.6G 82.4
+Mamba®-B [73] 280+20 99M 20.6G 82.9
+ViL-B 400+5 89M 17.9G 82.4
 Table 1: ImageNet-1K pre-training accuracy.
@@
 Models with “+” in their “Epochs” column pre-train on lower resolution followed by fine-tuning on 224x224 resolution for some epochs.
-ViL performs favorably against against an isotropic convolutional architecture (ConvNeXt) and vision adaptions of transformers (DeiT), RWKV [47] (VRWKV) and Mamba [23] (Vim, Mamba®).
+ViL performs favorably against an isotropic convolutional architecture (ConvNeXt) and vision adaptions of transformers (DeiT series), RWKV (VRWKV) and Mamba (Vim, Mamba®).
+Appendix Table 9 confirms these results on OOD and robustness evaluations of these classifiers.
 3.2 ADE20K Semantic Segmentation
@@
 Also here, ViL shows strong performances across the board, even outperforming DeiT-III-B despite the lower ImageNet-1K accuracy of ViL-B.
-For this benchmark, we train on 512x512 images and inference could be conducted on the original images (without resizing), which results in even larger images.
-Consequently, the number of patch tokens can become quite large.
-As sequence length increases, ViL would ideally use a recurrent or chunkwise formulation of the mLSTM layer.
-Shortly summarized, the mLSTM can operate in parallel, recurrent or chunked modes, each with distinct FLOPS and runtime characteristics.
-Given a sequence length $L$ and hidden dimension $D$, the trade-off between modes can be formulated as $\mathcal{O}(LCD+LD^{2})$ where $C$ denotes the chunksize of the chunked form.
-$C=1$ (recurrent mode) yields complexity $\mathcal{O}(LD^{2})$ and $C=L$ (parallel mode) gives $\mathcal{O}(L^{2}D)$ or $\mathcal{O}(L\frac{L}{2}D)$ when omitting the causally masked operations.
-Due to this trade-off and the current lack of an optimized hardware implementation for the mLSTM, it is difficult to create runtime or FLOPS estimates that are of practical relevance in this setting and therefore leave it to future work.
+The high resolution of the ADE20K segmentation task (512x512) results in a total of 1024 patch tokens where the quadratic complexity of self-attention is significantly more expensive than the linear complexity of the mLSTM, resulting in much fewer FLOPS for ViL.
+Additionally, the efficient alternating block design results in lower FLOPS than Mamba-based vision models (which also have linear complexity).
 Single-scale Multi-scale
-Model #Params mIoU ACC mIoU ACC
-DeiT-T [51] 10M 38.1 78.2 40.3 79.9
-DeiT-III-T [54] 10M 39.8 79.2 42.2 80.7
-Vim-T [67] 13M 41.0 - - -
-ViL-T 11M 41.2 80.2 43.1 81.3
-DeiT-S [51] 41M 43.1 80.7 45.2 81.8
-DeiT-III-S [54] 41M 45.2 81.5 46.3 82.3
-Vim-S [67] 46M 44.9 - - -
-Mamba®-S [58] 56M 45.3 - - -
-ViL-S 42M 46.3 82.0 47.9 82.9
-DeiT-B [51] 113M 45.8 82.1 47.0 82.9
-DeiT-III-B [54] 113M 47.5 82.6 49.0 83.3
-Mamba®-B [58] 132M 47.7 - - -
-ViL-B 115M 48.6 82.8 49.6 83.3
+Model #Params FLOPS mIoU ACC mIoU ACC
+DeiT-T 10M 10.4G 38.1 78.2 40.3 79.9
+DeiT-III-T 10M 10.4G 39.8 79.2 42.2 80.7
+Vim-T 13M 7.7G 41.0 - - -
+ViL-T 11M 6.6G 41.2 80.2 43.1 81.3
+DeiT-S 41M 31.7G 43.1 80.7 45.2 81.8
+DeiT-III-S 41M 31.7G 45.2 81.5 46.3 82.3
+Vim-S 46M 27.3G 44.9 - - -
+Mamba®-S 56M 27.6G 45.3 - - -
+ViL-S 42M 24.4G 46.3 82.0 47.9 82.9
+DeiT-B 113M 107.0G 45.8 82.1 47.0 82.9
+DeiT-III-B 113M 107.0G 47.5 82.6 49.0 83.3
+Mamba®-B 132M 102.8G 47.7 - - -
+ViL-B 115M 93.6G 48.6 82.8 49.6 83.3
 Table 2: Semantic segmentation results on ADE20K [81] using UperNet [75].
@@
 Models are trained for 160K updates with a batchsize of 16 on 512x512 resolution.
+We use a feature pyramid consisting of rescaled feature maps after the 4th, 6th, 8th and final block.
 Detailed hyperparameters are listed in Appendix Table 12.
+FLOPS are calculated only from the backbone at 512x512 resolution as all models use the same segmentation head.
 3.3 VTAB-1K Transfer Classification
+Model #Params FLOPS Natural Specialized Structured Average
+DeiT-T 6M 1.3G 69.2 82.0 53.3 65.2
+DeiT-III-T 6M 1.3G 71.9 82.6 55.2 67.1
+Vim-T 7M 1.5G 68.0 80.7 47.1 61.9
+ViL-T 6M 1.3G 73.6 83.4 56.1 68.3
+DeiT-S 22M 4.6G 73.3 83.8 53.2 67.1
+DeiT-III-S 22M 4.6G 75.0 83.2 52.3 67.2
+Vim-S 26M 5.3G 69.6 81.7 49.4 63.6
+ViL-S 23M 4.7G 75.3 84.3 58.3 70.0
+DeiT-B 86M 17.6G 76.5 85.2 55.7 69.6
+DeiT-III-B 86M 17.6G 77.6 84.8 56.6 70.3
+ViL-B 89M 17.9G 76.6 84.7 59.1 70.9
+Table 3: Transfer classification accuracies on the VTAB-1K [79] benchmark using ImageNet-1K pre-trained models.
+VTAB-1K consists of 19 datasets split into 7 natural, 4 specialized and 8 structured datasets.
+We show averages per category and the average accuracy over all 19 datasets (Appendix Table 8 lists all individual accuracies).
+ViL shows strong generalization performance, outperforming heavily optimized ViT protocols and Vim on the full VTAB-1K benchmark.
+ViL performs exceptionally well on the structured category.
+We tune the learning rate for each model and dataset on the validation set and report the average testset accuracy over 5 seeds.
+Appendix Table 11 lists further hyperparameters.
 Table 3 shows transfer classification results for ImageNet-1K pre-trained models on the VTAB-1K [79] benchmark.
-VTAB-1K consists of 19 datasets split into 7 natural datasets (such as CIFAR100 [38], Caltech101 [20] or SUN397 [60]), 4 specialized datasets (medical imaging [57, 34] and remote sensing [28, 11]) and 8 structured datasets (with tasks such as object counting [33] or binned depth estimation [21]).
+VTAB-1K consists of 19 datasets split into 7 natural datasets (such as CIFAR100 [45] or Caltech101 [24]), 4 specialized datasets (medical imaging [71, 43] and remote sensing [35, 14]) and 8 structured datasets (with tasks such as object counting [42] or binned depth estimation [26]).
 We follow common practices and tune the learning rate per model and dataset on the validation set followed by training with the best learning rate on the union of train and validation set.
@@
 ViL performs particularly well on the structured datasets where ViL-B outperforms DeiT-III-B despite ViL-B having lower ImageNet-1K accuracy.
-Model #Params FLOPS Natural Specialized Structured Average
-DeiT-T [51] 6M 1.3G 69.2 82.0 53.3 65.2
-DeiT-III-T (reimpl.) 6M 1.3G 71.9 82.6 55.2 67.1
-Vim-T [67] 7M 1.5G 68.0 80.7 47.1 61.9
-ViL-T 6M 1.5G 73.6 83.4 56.1 68.3
-DeiT-S [51] 22M 4.6G 73.3 83.8 53.2 67.1
-DeiT-III-S [54] 22M 4.6G 75.0 83.2 52.3 67.2
-Vim-S [67] 26M 5.3G 69.6 81.7 49.4 63.6
-ViL-S 23M 5.1G 75.3 84.3 58.3 70.0
-DeiT-B [51] 86M 17.6G 76.5 85.2 55.7 69.6
-DeiT-III-B [54] 86M 17.6G 77.6 84.8 56.6 70.3
-ViL-B 89M 18.6G 76.6 84.7 59.1 70.9
-Table 3: Transfer classification accuracies on the VTAB-1K [64] benchmark using ImageNet-1K pre-trained models.
-VTAB-1K consists of 19 datasets split into 7 natural, 4 specialized and 8 structured datasets.
-We show averages per category and the average accuracy over all 19 datasets (Appendix Table 10 lists all individual accuracies).
-ViL shows strong transfer classification performance, outperforming heavily optimized ViT protocols and Vim on the full VTAB-1K benchmark.
-Notably, ViL performs exceptionally well on the structured category.
-We tune the learning rate for each model and dataset on the validation set and report the average performance on the test set over 5 seeds.
-Appendix Table 13 lists further hyperparameters.
-3.4 Block Design
-We study different ways to design ViL blocks in Table 4.
-The plain, uni-directional xLSTM block fails to reach competitive performances since the autoregressive nature of xLSTM is not suited for image classification.
-Traversing blocks in a bi-directional manner, i.e., introducing a second mLSTM layer in each block that traverses the sequence backwards (akin to Vim [67]), improves performance, but also requires more parameters and FLOPS.
-Sharing the parameters of the forward and backward mLSTM makes the model more parameter efficient, but still requires more compute and overloads these parameters which leads to performance drops.
-Using alternating blocks enhances performance while remaining compute and parameter efficient.
-We also explore quad-directional designs (similar to [40]), which refers to traversing the sequence row-wise (in both directions) and additionally column-wise (in both directions).
-Bi-directional traverses the sequence only row-wise (in both directions).
-Figure 3 visualizes the different traversal paths.
-Due to the increased cost of bi-directional and quad-directional blocks, this study was conducted in a heavily reduced setting.
-We train on a subset of ImageNet-1K that contains only samples from 100 classes for 400 epochs at 128x128 resolution.
-This was particularly necessary because our quad-directional implementation is not compatible with torch.compile (a generic speed optimization method from PyTorch [45]), which leads to longer runtimes, as indicated last column of Table 4.
-Due to this technical limitation, we choose the alternating bi-directional block as our core design.
-Direction Alternating Param. Sharing ACC #Param FLOPS Runtime
-Uni-dir. (xLSTM) 72.8 6.1M 1.0x 1.0x
-Bi-dir. 74.8 6.5M 1.4x 1.9x
-Bi-dir. 73.3 6.1M 1.4x 1.9x
-Bi-dir. (ViL) 74.8 6.1M 1.0x 1.0x
-Quad-dir. 76.3 7.3M 2.3x 7.5x
-Quad-dir. 72.8 6.1M 2.3x 7.5x
-Quad-dir. 75.6 6.1M 1.0x 1.9x
-Table 4: Traversing the sequence of patch tokens in multiple directions yields better performance than the plain, uni-directional xLSTM block.
-However, using multiple directions per block also requires more compute and parameters.
-Sharing parameters between directions matches #Params but overloads them, leading to performance drops.
-Using multi-directional designs but alternating between them in consecutive blocks, matches #Params and FLOPS while achieving good performances.
-While quad-directional designs yields the best performances, they comes with drastically increased runtime.
-As our current quad-directional implementation is not compatible with torch.compile, it suffers from increased runtime even when alternating between directions.
-Since quad-directional designs are compute intensive, this study was conducted by training a ViL-T on an ImageNet-1K subset of 100 classes using resolution 128.
-“Param.
-Sharing” shares the parameters of the two (bidirectional) or four (quaddirectional) mLSTM layers in each block.
-Figure 3: Visualization of traversal paths. Bi-directional uses the first two traversal paths. Quad-directional uses all of the above. Squares represent individual patch tokens.
+4 Ablation Studies
+We ablate various design choices of ViL by training ViL-T models for 100 epochs on ImageNet-1K in 224x224 resolution, other hyperparameters follow the ones from Section 3 (see also Appendix B.3).
+We then report the validation accuracy on ImageNet-1K and fine-tune the model on ADE20K to ensure that design choices are not overfitted to classification.
+We also use a reduced segmentation pipeline where we use a linear segmentation head and train for 40K updates using a batch size of 16 (other hyperparameters follow Appendix 12).
+4.1 Architectural Design
+We consider various architecture design choices in Table 4.
+Directions IN1K ADE20K
+Uni-dir. 72.2 28.6
+Bi-dir. 73.7 31.7
+Quad-dir. 73.8 33.1
+Oct-dir. 73.5 32.4
+Convolution IN1K ADE20K
+None 72.3 29.2
+Causal-Conv1D 72.8 27.8
+Conv1D 72.8 28.4
+Conv2D 73.7 31.7
+Pos. Embed. IN1K ADE20K
+73.7 31.0
+73.7 31.7
+Concurrency IN1K ADE20K
+Sequential 73.7 31.7
+Parallel 73.0 30.6
+Table 4: Architecture design ablation studies. Default settings
+Figure 4: Uni-directional, bi-directional, quad-directional and oct-directional traversal paths.
+Squares represent individual patch tokens.
+Traversal starts at the circle and goes in direction of the arrow, if no further patches are in a row/column, the traversal continues in the next row/column as indicated by the dashed line.
+(a) Traversal Directions
+Traversing the sequence in at least two directions greatly improves performance due to the non-causal 2D structure of images.
+Adding column-wise traversal directions (Quad-dir.) could even further improve semantic segmentation performance.
+Additionally using 4 instead of 2 starting positions (Oct-dir.) shows no benefit.
+Note that all variants have the same amount of FLOPS due to sequential application of different directions.
+Directions are visualized in Figure 4.
+We use “Bi-dir.” for our final models due to current technical limitations which would slow down training on more than 2 directions.
+This limitation comes from the current lack of optimized hardware implementations of the mLSTM (e.g., CUDA kernels) where we instead rely on torch.compile, a generic speed optimization method from PyTorch [56], to optimize computations.
+Our implementation of quad- and oct-directional traversals is not compatible with torch.compile, which results in approximately double the runtime.
+We therefore train all models from Section 3 with “Bi-dir.”.
+Note that this is merely a technical limitation, not a methodical one and the ablation study suggest that future ViL models could be even better using a quad-directional design.
+(b) QK Convolution
+The mLSTM block design uses a causal 1D convolution to aggregate local context to improve storage/retrieval to/from the cell state ${\bm{C}}$.
+This is done by applying a convolution layer to ${\bm{X}}$ before projecting it to ${\bm{Q}}$ with ${\bm{W}}_{q}$ and ${\bm{K}}$ with ${\bm{W}}_{k}$ respectively.
+The convolution is shared for ${\bm{Q}}$ and ${\bm{K}}$.
+The causal 1D structure of the convolution from the original mLSTM [5] is necessary due to the causal 1D structure of language modeling.
+However, as images are neither causal nor 1D structures, we replace the causal 1D convolution with a 2D convolution (with kernel size 3).
+This allows the mLSTM to make better storage/retrieval decisions through the added local context.
+(c) Positional Embedding
+ViTs require positional embedding to tell the model where each patch is located in the image, suffering heavy performance losses if the position is not required [22, 15].
+The mLSTM is an autoregressive model, which makes it optional to add positional embeddings as it can recognize the position of the current patch based on how many patches have been processed.
+However, the ablation shows that it is nevertheless beneficial to provide this information explicitly as it improves segmentation results without hurting classification performance.
+(d) Sequential vs. Parallel
+Related architectures use a parallel design where a sequence is processed from multiple directions in a single block [82, 23].
+We investigate a similar design where we apply both directions in parallel instead of sequentially.
+To keep parameters and FLOPS constant, we apply the directions akin to parallel transformer blocks [72] while halving the depth.
+${\bm{Y}}={\bm{X}}+\text{Block}_{\theta}({\bm{X}})+\text{Flip}(\text{Block}_{\phi}(\text{Flip}({\bm{X}})))$ (15)
 4.2 Classification Design
-In order to perform classification with a ViT, the sequence of tokens is pooled into a single token, which is then used as input to a classification head.
-The most common methods to do this pooling are (i) adding a learnable [CLS] token at the start of the sequence or (ii) averaging all patch tokens to produce an [AVG] token.
-Whether to use the [CLS] or [AVG] token is typically a hyperparameter, where both variants achieve comparable performances.
-On the contrary, autoregressive models often require specialized classification designs.
+In order to perform classification from a sequence of tokens, it is common to aggregate information from the whole sequence, which is then used as input to a classification head.
+The most common methods to do this aggregation are (i) adding a learnable [CLS] token to the input sequence or (ii) averaging all patch tokens to produce an [AVG] token.
+In ViTs, whether to use the [CLS] or [AVG] token is typically a hyperparameter, where both variants achieve comparable performances.
+On the contrary, other sequence models models often require specialized classification designs.
 For example, Vim [82] requires the [CLS] token to be in the middle of the sequence, suffering heavy performance losses if other classification designs, e.g., an [AVG] token or two [CLS] tokens at start and end of the sequence, are employed.
-Due to its autoregressive nature, we explore different classification designs for ViL in Table 5.
-[AVG] averages all patch tokens, “Middle Patch” uses the middle patch token, “Middle [CLS]” uses a [CLS] token in the middle of the sequence, “Bilateral [AVG]” uses the average of the first and last patch token.
-We find ViL to be relatively robust the classification design where all performances that use a single token are within 0.6%.
-We only add a [CLS] token if it is used for classification, hence the best design (“Bilateral Concat”) does not use a [CLS] token.
-Design Dim ACC
-[AVG] 192 75.4
-Middle Patch 192 75.9
-Middle [CLS] 192 75.3
-Bilateral [AVG] 192 75.9
-Bilateral Concat 384 76.4
-Table 5: How to pool the sequence of tokens for classification?
-ViL is relatively robust to the classification design.
-Using the middle patch token and averaging the first and last token (“Bilateral [AVG]”) work similarly well.
-The good performance of middle patch could be caused by the bias in ImageNet-1K that objects are typically in the center of the image.
-We therefore choose a bilateral design to avoid exploiting this bias.
-Additionally, concatenating the first and last token (“Bilateral Concat”) instead of averaging improves performance and is used as final design for ViL.
-We conduct this study by training a ViL-T on ImageNet-1K for 400 epochs.
-3.6 Vision Adjustments of xLSTM
-The xLSTM [5] architecture was designed for language modeling and therefore includes some design choices that are specific to language modeling.
-One of these design choices is a causal 1D convolution (with kernel size 4) to encode local context into the query and key input of the mLSTM layer.
-As images are neither causal nor 1D structures, we replace the causal 1D convolution with a 2D convolution (with kernel size 3).
-When training a ViL-T for 400 epochs, this improves ImageNet-1K accuracy from 76.1% to 76.7%.
-Another language specific design choice is the absence of bias parameters in projection and layernorm [4] layers which improves training stability for larger models [12, 16].
-As vision models are typically much smaller than language models, ViTs commonly use all bias parameters.
-When adding bias parameters, a ViL-T trained for 400 epochs improves from 76.7% to 77.1% on ImageNet-1K.
+We explore different classification designs for ViL in Table 5.
+(a) We choose concatenating the first and last patch as aggregation method due to its strong classification performance.
+As our final models also perform well in semantic segmentation (see Table 2), we do not retrain models with [AVG] aggregation even though the ablation suggests that this could boost performance even further for segmentation tasks.
+(b) Adding learnable [CLS] tokens show no benefit.
+Therefore, we do not use any [CLS] tokens for ViL.
+Aggregation IN1K ADE20K
+Bilateral Mean 73.0 31.5
+Bilateral Concat 73.7 31.7
+[[[[AVG]]]] 72.6 32.8
+Center [AVG] 72.4 32.1
+Aggregation IN1K
+Concat Bilateral Patches 73.7
+Mid [CLS] 71.8
+Bilateral [CLS] 73.5
+Mid + Bilateral [CLS] 73.0
+Table 5: Classification design.
+(a) ViL aggregates classification information well in the first and the last patches (bilateral), leading to good classification performance if the first and last patches are averaged or concatenated.
+Averaging all patches ([AVG]) or the 4 center patches (Center [AVG]) results in strong segmentation performances but lackluster classification performances.
+(b) Adding learnable [CLS] tokens to the start and end of the input sequence (Bilateral [CLS]) offers no benefit over simply using the first and the last patch.
+Incorporating a [CLS] token in the middle of the sequence, akin to Vim [82], does not improve performance.
+Default settings
+5 Limitations and Future Work
+The biggest limitation of ViL is the current lack of an optimized hardware implementation of the mLSTM, which results in longer runtimes than ViTs, which have multiple optimized hardware implementations [18, 17].
+This makes a runtime/throughput analysis of models, a vital metric to judge practicability, difficult as the practical relevance of inefficient implementations is quite low.
+As a proxy, we report FLOP counts, where ViL is comparable to ViT on low-resolution tasks and far better than ViT on high-resolution tasks due to its linear complexity.
+While FLOPS are far from an optimal proxy for runtime/throughput, they suggest that ViL can be much faster than ViT on high-resolution tasks once an optimized hardware implementation exists.
+Note that ViL is already faster than Vim (see Appendix A.1) despite its optimized hardware implementation.
+This limitation snowballs in multiple other directions.
+For example, scaling model size further, tuning hyperparameters, training on larger datasets, exploring self-supervised pre-training or investigating hierarchical architectures are all interesting avenues for future work that are currently quite costly due to the lack of an optimized hardware implementation.
+Please note that this is merely a technical limitation, not a methodical one as the mLSTM is heavily parallelizable.
+However, implementing fast compute kernels in CUDA [54] or Triton [63] is highly non-trivial as it requires expert hardware architecture knowledge, advanced implementation skills and potentially multiple development cycles to iron out numerical inaccuracies or instabilities.
+However, the results of recent linear attention mechanisms show impressive FLOPS utilization (e.g., [78]).
+As the mLSTM can be parallelized with similar techniques it is only a matter of time that the mLSTM achieves a similar FLOPS utilization, which will make the mLSTM faster than transformers once an efficient hardware implementation is available.
+Additionally, we made a significant effort to make our architecture as efficient as possible, using the tools that are currently available to us.
+Notably, our architecture is already much faster (up to 70%) than Vim [82] despite Vim using a custom CUDA kernel, as shown in Appendix A.1.
+For reference, in language modeling, Mamba is roughly on-par with transformers in terms of speed and 4x faster than than the xLSTM (as mentioned in [5]), again, due to the current lack of efficient hardware implementation of the mLSTM.
+These considerations further underline the potential of our simple and efficient design for vision applications.
+6 Related Work
+Generic Vision Backbones.
+The inductive bias of CNNs [25, 47] has demonstrated ground-breaking advancements in computer vision [46] in the early deep learning days.
+Features of CNNs have been found to learn generic visual features that can be used for a variety of tasks [21].
+Subsequently, countless works improved various aspects such as architectures [60, 33, 41, 61, 50] or pre-training strategy [20, 53, 80, 27, 12, 28].
+Sequence Models in Vision.
+The introduction of transformers [70] demonstrated exceptional scalability in language processing, which motivated the vision community to explore transformers also in computer vision [11, 16] but was applied on pixels or small patches which inhibited large costs due to the quadratic complexity of self-attention.
+This restriction was alleviated by the seminal work Vision Transformers (ViTs) [22] by using larger patches to aggregate local information and reduce training costs.
+Similar to CNNs, lots of work improved on the ViT architecture by refining training procedures [64, 65, 67, 9, 4, 76, 34].
+The recent advancement of autoregressive models in language processing [30, 58] has also gathered interest in the vision community [82, 23] due to the linear scaling property which allows applications to high-resolution tasks such as medical imaging [51] or video understanding [48].
 7 Conclusion
@@
 ViL processes a sequence of patch tokens in alternating fashion.
-Odd blocks process image patches row-wise from top left to bottom right and even blocks go from bottom right to top left.
-Our new architecture outperforms SSM-based vision architectures and also optimized ViT models on ImageNet-1K classification, VTAB-1K transfer classification and ADE20K semantic segmentation.
+Odd blocks process image patches row-wise from top left to bottom right and even blocks go row-wise from bottom right to top left.
+Our new architecture outperforms SSM-based vision architectures, other autoregressive vision architectures and also optimized ViT models on ImageNet-1K classification, VTAB-1K transfer classification and ADE20K semantic segmentation.
 Remarkably, ViL is able to outperform ViT training pipelines, which are the result of years of hyperparameter tuning and transformer improvements.
 In the future, we see potential in applying ViL when high-resolution images are needed for optimal performance, such as semantic segmentation or medical imaging.
-In these settings, transformers suffer from high computational costs due to the quadratic complexity of self-attention, where ViL can use a chunked form to trade-off between its parallel form (quadratic complexity) and recurrent form (linear complexity).
-Additionally, improving pre-training schemes (e.g.
-via self-supervised learning), exploring better hyperparameter settings or investigating hierarchical architectures are promising future directions that could improve ViL even further.
+In these settings, transformers suffer from high computational costs due to the quadratic complexity of self-attention, where the linear complexity of ViL allows compute efficient processing of long sequences.
+Additionally, improving pre-training schemes (e.g., via self-supervised learning), exploring better hyperparameter settings or investigating hierarchical architectures are promising future directions.
 Acknowledgments
@@
 Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.
-Appendix A Changelog
-A.1 Version 1
-In the first version of this paper, we trained models with a causal 1D convolution in the mLSTM layer (instead of a 2D convolution), “Bilateral [AVG]” pooling (instead of “Bilateral Concat”) and a constant resolution of 224 (instead of pre-training at 192 followed by a short fine-tuning on resolution 224).
-This version of the architecture is also used to conduct the studies in Table 4 and Table 5.
-For completeness, we report the accuracies of these models in Table 6.
-In the first version, we trained slightly deeper models (ViL-T+ with 29 blocks and ViL-S+ with 26 blocks) to match the parameter count of Vim-T and Vim-S.
-However, as the models with a regular depth of 24 were already better than Vim with less parameters, we omit them in future versions.
-Method #Params ACC
-ViL-T 6M 77.3
-ViL-T+ 7M 78.1
-ViL-S 23M 80.7
-ViL-S+ 26M 80.9
-ViL-B 88M 81.6
-Model #Params ACC
-Long-sequence fine-tuning
-ViL-T+ 7M 80.0
-ViL-S+ 26M 82.2
-ViL-B 88M 82.7
-Table 6: ImageNet-1K accuracies of models trained in the first version of this paper.
-A.2 Version 2
-The current version of the paper improves performance by (i) replacing the causal 1D convolution with a 2D convolution (with kernel size 3 and padding 1) (ii) using “Bilateral Concat” instead of “Bilateral [AVG]” pooling (iii) using bias parameters in projection and LayerNorm [4] layers (iv) pre-training models on 192x192 followed by a short fine-tuning on 224x224 resolution.
-Appendix B Extended Results
-B.1 Runtime Comparison of ViL vs Vim
-We compare the time it takes to train ViL and Vim [67] for a single ImageNet-1K epoch in Table 7.
-We follow the scaling procedure of ViTs, using 192 (T), 384 (S), 768 (B), 1024 (L) as latent dimension where the large scale additionally doubles the number of blocks.
+Appendix A Extended Results
+A.1 Runtime Comparison of ViL vs Vim
+We compare the runtime to train ViL and Vim [82] for 10 ImageNet-1K epochs in Table 6.
+We follow the scaling procedure of ViTs, using 192 (T), 384 (S), 768 (B), 1024 (L) as hidden dimension where the (L)arge scale doubles the number of blocks.
 Model Optimization (T)iny (S)mall (B)ase (L)arge
@@
 Table 6: Runtime comparisons between Vim [82] and ViL.
-ViL is up to 69% faster despite the current lack of a custom CUDA kernel.
+ViL is up to 69% faster despite the current lack of a optimized hardware implementation.
 As mLSTM (and ViL) can be parallelized analogous to FlashAttention [18, 17] via custom hardware optimizations, ViL will become even faster in the future.
 Runtimes denote the training time for 10 ImageNet-1K epochs and are extrapolated from short benchmark runs on a single A100-80GB-PCIe using float16 precision and 224x224 images.
-B.2 Impact of Longer Training
+A.2 Impact of Longer Training
 We investigate the impact of training for a longer duration in Table 7.
-We also attempted to train larger models for longer but did not observe similar performance gains.
-One could probably change that by tuning hyperparameters for an 800 epoch training schedule, which we omit due to the large computational costs associated with training larger models.
 Model Epochs IN-1K ACC VTAB-1K ADE20K mIoU
@@
 ViL-T 800 78.3 68.3 41.2
-Table 8: Performance comparison of tiny models when training for 400 and 800 epochs. ADE20K mIoU uses single-scale evaluation. All settings follow the ones used in the main paper.
-B.3 Long-sequence fine-tuning
-Table 9 shows results for the long-sequence fine-tuning setting of Vim [67].
-Method #Params ACC
-DeiT-III-T (reimpl.) 6M 78.3
-Vim-T [67] 7M 78.3
-ViL-T 6M 80.1
-DeiT-S [51] 22M 81.4
-DeiT-III-S [54] 22M 83.2
-Vim-S [67] 26M 81.6
-ViL-S 23M 82.7
-DeiT-B [51] 86M 83.2
-DeiT-III-B [54] 86M 84.5
-ViL-B 89M 83.3
-Table 9: Fine-tuning ImageNet-1K pre-trained models by using a 50% overlap when creating patches.
-This setting increases the number of patch tokens from 196 to 729.
-ViL easily outperforms Vim but falls short of ViTs on larger models.
-Note that ViL FLOPS are calculated using the parallel form of the mLSTM, which scales quadratically with the sequence length.
-Using the chunkwise form instead of the parallel form can reduce FLOPS for longer sequences (see Section C.2).
-B.4 VTAB-1K Individual Dataset Results
-Table 10 presents accuracies for each individual dataset of the VTAB-1K [64] benchmark.
+Table 7: Performance comparison of tiny models trained for 400 and 800 epochs. ADE20K mIoU uses single-scale evaluation. All settings follow the ones used in the main paper.
+A.3 VTAB-1K Individual Dataset Results
+Table 8 presents accuracies for each individual dataset of the VTAB-1K benchmark.
 Natural Specialized Structured
@@
 sNORB-Ele
-DeiT-T [51] 47.7 86.4 63.7 85.6 87.0 78.4 35.3 83.0 93.4 80.9 70.7 71.7 60.3 43.1 78.5 67.9 41.6 30.6 32.7
-DeiT-III-T (reimpl.) 52.3 90.1 62.7 88.8 87.5 83.7 37.9 83.2 93.1 81.1 72.9 76.6 60.8 44.9 79.1 67.5 48.1 31.0 33.3
-Vim-T [67] 46.7 86.3 60.7 84.0 88.8 76.1 33.7 82.2 92.9 75.2 72.6 59.8 49.9 39.3 78.2 51.2 43.9 26.9 27.2
+DeiT-T 47.7 86.4 63.7 85.6 87.0 78.4 35.3 83.0 93.4 80.9 70.7 71.7 60.3 43.1 78.5 67.9 41.6 30.6 32.7
+DeiT-III-T 52.3 90.1 62.7 88.8 87.5 83.7 37.9 83.2 93.1 81.1 72.9 76.6 60.8 44.9 79.1 67.5 48.1 31.0 33.3
+Vim-T 46.7 86.3 60.7 84.0 88.8 76.1 33.7 82.2 92.9 75.2 72.6 59.8 49.9 39.3 78.2 51.2 43.9 26.9 27.2
 ViL-T 54.2 90.2 67.4 90.7 89.9 81.6 41.1 83.4 94.2 82.7 73.1 80.7 61.8 49.4 81.3 57.8 51.8 31.4 34.8
-DeiT-S [51] 57.0 88.9 68.2 90.9 90.8 75.4 42.1 83.3 94.0 83.8 74.0 74.6 58.3 45.6 78.2 61.9 47.9 27.1 31.9
-DeiT-III-S [54] 58.8 88.6 67.5 90.9 91.7 84.4 43.3 84.4 92.6 82.5 73.5 76.5 57.9 46.2 78.9 58.3 49.7 23.7 27.5
-Vim-S [67] 53.0 87.2 64.1 86.8 90.3 65.8 39.7 82.4 93.4 78.0 73.1 63.1 53.2 42.3 78.2 54.1 47.6 27.1 29.3
+DeiT-S 57.0 88.9 68.2 90.9 90.8 75.4 42.1 83.3 94.0 83.8 74.0 74.6 58.3 45.6 78.2 61.9 47.9 27.1 31.9
+DeiT-III-S 58.8 88.6 67.5 90.9 91.7 84.4 43.3 84.4 92.6 82.5 73.5 76.5 57.9 46.2 78.9 58.3 49.7 23.7 27.5
+Vim-S 53.0 87.2 64.1 86.8 90.3 65.8 39.7 82.4 93.4 78.0 73.1 63.1 53.2 42.3 78.2 54.1 47.6 27.1 29.3
 ViL-S 61.4 89.6 69.2 92.8 91.7 78.7 43.8 85.5 93.9 84.4 73.5 84.0 63.4 51.3 83.3 61.0 55.4 32.4 35.5
-DeiT-B [51] 61.8 89.8 67.5 93.7 92.6 84.4 45.6 85.3 95.1 86.3 74.2 77.7 59.9 47.2 81.7 61.7 51.4 30.0 36.2
-DeiT-III-B [54] 62.9 89.6 69.6 93.7 93.2 87.0 47.1 85.8 94.1 85.6 73.7 80.5 61.4 48.4 80.9 64.4 55.1 30.2 31.8
+DeiT-B 61.8 89.8 67.5 93.7 92.6 84.4 45.6 85.3 95.1 86.3 74.2 77.7 59.9 47.2 81.7 61.7 51.4 30.0 36.2
+DeiT-III-B 62.9 89.6 69.6 93.7 93.2 87.0 47.1 85.8 94.1 85.6 73.7 80.5 61.4 48.4 80.9 64.4 55.1 30.2 31.8
 ViL-B 64.3 90.0 71.1 93.4 91.4 79.6 46.6 84.7 94.3 85.4 74.4 83.7 62.1 52.7 81.0 63.1 57.6 32.6 39.9
 Table 8: Results on all datasets of the VTAB-1K [79] benchmark.
-B.5 Robustness
+A.4 Robustness and Domain Generalization
 Table 9 presents robustness and OOD evaluations of ImageNet-1K pre-trained classifiers.
 Model IN-C ($\downarrow$) IN-A ($\uparrow$) IN-R ($\uparrow$) Sketch ($\uparrow$) Validation ($\uparrow$)
-DeiT-T [51] 69.7 7.6 32.7 19.9 72.2
+DeiT-T 69.7 7.6 32.7 19.9 72.2
 DeiT-III-T 65.0 11.7 39.4 27.4 76.2
-Vim-T [67] 61.8 9.6 38.8 26.9 76.1
+Vim-T 61.8 9.6 38.8 26.9 76.1
 ViL-T 59.6 15.2 42.2 30.0 78.3
-DeiT-S [51] 54.4 19.6 41.9 29.1 79.8
-DeiT-III-S [54] 50.1 23.2 46.6 35.4 81.4
-Vim-S [67] 51.5 19.7 44.8 32.5 80.5
+DeiT-S 54.4 19.6 41.9 29.1 79.8
+DeiT-III-S 50.1 23.2 46.6 35.4 81.4
+Vim-S 51.5 19.7 44.8 32.5 80.5
 ViL-S 50.6 23.8 47.9 35.2 81.5
-DeiT-B [51] 48.6 27.9 44.6 32.0 81.8
-DeiT-III-B [54] 42.7 36.5 54.1 41.1 83.8
+DeiT-B 48.6 27.9 44.6 32.0 81.8
+DeiT-III-B 42.7 36.5 54.1 41.1 83.8
 ViL-B 45.3 30.9 51.9 39.0 82.4
@@
 For ImageNet-C, we report the mean corruption error [36] with AlexNet [46] as baseline.
-Appendix C Implementation Details
-C.1 Hardware
-We train models on a mix of custom hardware servers (mainly A100 and A40 GPUs) and public research clusters equiped with either 8xA100 or 4xA100 nodes.
+Appendix B Implementation Details
+B.1 Hardware
+We train models on servers with either 8xA100 or 4xA100 nodes.
 We estimate the total number of A100 GPU-hours used for this project to be 38K hours. This estimate includes initial exploration, method development, analysis and evaluations.
-C.2 FLOPS Calculation
-We calculate the FLOPS with the fvcore library [^1] https://github.com/facebookresearch/fvcore.
-For ViL, we report the FLOPS of the mLSTM parallel form.
-Given sequence length $L$ and hidden dimension $D$, it has complexity $\mathcal{O}(L\frac{L}{2}D)$ where the $\frac{L}{2}$ stems from the causal masking.
-We report FLOPS for a complexity of $\mathcal{O}(L\frac{L}{2}D)$ even though our current implementation of the mLSTM has complexity $\mathcal{O}(L^{2}D)$ (i.e., entries of the $QK$ matrix that are later masked are still calculated) which we justify by the fact that FlashAttention-2 [14] is approximately 1.7x faster with a causal mask than without.
-Therefore, an optimized hardware implementation of the mLSTM could also omit the calculation of values that are not needed.
+B.2 FLOPS Calculation
+We use the fvcorehttps://github.com/facebookresearch/fvcore library to count FLOPS and report FLOPS of the mLSTM chunkwise form as described in Section 2.2.
+For the parallel parts, we report FLOPS for a complexity of $\mathcal{O}\big{(}(\frac{S}{2}+1)Sd\big{)}$ because the upper triangular entries of the $\mathbf{QK}$ matrix do not need to be calculated due to the causal structure.
+We justify this by the fact that FlashAttention-2 [17] is approximately 1.7x faster with a causal mask than without.
+Therefore, an optimized hardware implementation of the mLSTM could also omit the calculation of the upper triangular part of $\mathbf{QK}$.
 As Vim [82] does not report FLOPS and their model makes use of CUDA kernels (which are not counted as FLOPS by fvcore), we replace all calls to CUDA kernels with their reference PyTorch implementation and count the FLOPS with fvcore.
-For the total pre-training compute in Figure 2, we consider an efficient implementation of stochastic depth [32] which omits the calculation of a dropped block instead of masking it.
+For the total pre-training compute in Figure 3, we consider an efficient implementation of stochastic depth [40, 68] which omits the calculation of a dropped block instead of masking it.
 Therefore, we change the implementation of ViT [22] to use our efficient stochastic depth implementation.
-Vim does not use stochastic depth for training as their models are relatively small.
-C.3 ViL Hyperparameters
+Vim does not use stochastic depth for training as they only train tiny and small models.
+B.3 ViL Hyperparameters
 Table 10 shows detailed hyperparameters used to train ViL models.
@@
    Pooling Bilateral Concat
-Stochastic depth [32]
+Stochastic depth
    Peak rate 0 (T), 0.05 (S), 0.2 (B)
    Layer-wise Decay
-Optimizer AdamW [42, 36]
+Optimizer AdamW
    Base Learning rate 1e-3 $\rightarrow$ 1e-5
-   Linear LR Scaling Divisor [22] 1024
+   Linear LR Scaling Divisor 1024
    Weight decay 0.05
@@
    Gradient Norm Clip 1.0
-Precision mixed bfloat16 [35]
+Precision mixed bfloat16
    Backend torch.autocast
@@
    End LR 1e-6
-Label smoothing [49]
+Label smoothing
 Train Data Augmentation
@@
    RandomHorizontalFlip $p=0.5$
-   3-Augment [54]
+   3-Augment
     Gaussian Blur $\sigma$ [0.1, 2.0]
@@
    Normalize ImageNet-1K statistics
-   Mixup [65] $\alpha$ 0.8
-   Cutmix [63] $\alpha$ 1.0
+   Mixup $\alpha$ 0.8
+   Cutmix $\alpha$ 1.0
 Test Data Augmentation
@@
 We follow the best setting from DeiT-III [67] and pre-train on 192 resolution followed by a short fine-tuning on 224 resolution (indicated by $\rightarrow$).
-C.4 Fine-tuning on VTAB-1K
+B.4 Fine-tuning on VTAB-1K
 For fine-tuning models on VTAB-1K we provide the hyperparameters in Table 11.
@@
 Seeds 5
-Optimizer AdamW [42, 36]
+Optimizer AdamW
    Learning rate [1e-3, 7.5e-4, 5.0e-4, 2.5e-4, 1.0e-4]
-   Layer-wise lr decay [13] 0.65*
+   Layer-wise lr deca 0.65*
    Weight decay 0.05
@@
    Warmup epochs 5
-Precision mixed bfloat16 [35]
+Precision mixed bfloat16
    Backend torch.autocast
@@
 *For Vim and ViL we group two consecutive blocks for the layer-wise lr decay similar to how ViT considers a pair of attention and MLP block as a single “layer” for the decay.
-C.5 ADE20K Semantic Segmentation Fine-tuning
+B.5 ADE20K Semantic Segmentation Fine-tuning
 We fine-tune models on ADE20K [81] using an UperNet [75] head.
@@
     Dimension 192 (T), 384 (S, B)
-Stochastic depth [32]
+Stochastic depth
    Peak rate 0 (T), 0.05 (S), 0.1 (B)
    Layer-wise Decay
-Optimizer AdamW [42, 36]
+Optimizer AdamW
    Learning rate 5e-4
-   Linear LR Scaling Divisor [22] 16
-   Layer-wise lr decay [13] 0.65*
+   Linear LR Scaling Divisor 16
+   Layer-wise lr decay 0.65*
    Weight decay 0.05
@@
 Table 12: Hyperparameters for fine-tuning on VTAB-1K. *For ViL we group two consecutive blocks into one similar to how a ViT block consists of a pair of attention and MLP block.
-C.6 DeiT-III Reimplementation Hyperparameters
+B.6 DeiT-III Reimplementation Hyperparameters
 Table 10 shows detailed hyperparameters used to train DeiT-III-T (reimpl.) from Table 1.
+Our reimplementation easily outperforms older baselines like DeiT-II-T (+2.7% ImageNet-1K accuracy) and is approximately even with the original on ADE20K (40.1 vs 39.8 on mIoU single-scale, 41.8 vs 42.2 mIoU multi-scale).
 Parameter Value
@@
    Pooling [CLS]
-Stochastic depth [32]
-Layerscale [52] 1e-4
-Optimizer AdamW [42, 36]
+Stochastic depth
+Layerscale 1e-4
+Optimizer AdamW
    Base Learning rate 1e-3 $\rightarrow$ 1e-5
-   Linear LR Scaling Divisor [22] 1024
+   Linear LR Scaling Divisor 1024
    Weight decay 0.05
@@
    Gradient Norm Clip
-Precision mixed bfloat16 [35]
+Precision mixed bfloat16
    Backend torch.autocast
@@
    End LR 1e-6
-Label smoothing [49]
+Label smoothing
 Train Data Augmentation
@@
    RandomHorizontalFlip $p=0.5$
-   3-Augment [54]
+   3-Augment
     Gaussian Blur $\sigma$ [0.1, 2.0]
@@
    Normalize ImageNet-1K statistics
-   Mixup [65] $\alpha$ 0.8
-   Cutmix [63] $\alpha$ 1.0
+   Mixup $\alpha$ 0.8
+   Cutmix $\alpha$ 1.0
 Test Data Augmentation